### v1

# UC-MOA: Utility-Conditioned Multi-Objective Alignment for Distributional Pareto-Optimality

Abstract
Reinforcement Learning from Human Feedback (RLHF) has become a cornerstone for aligning large language models (LLMs) with human values. However, existing approaches struggle to capture the multi-dimensional, distributional nuances of human preferences. Methods such as RiC that directly inject raw reward values into prompts face significant numerical sensitivity issues—for instance, LLMs may fail to distinguish between 9.11 and 9.8—while alternatives like MORLHF, Rewarded Soups, and MODPO incur high computational costs by training multiple models. In this work, we introduce Utility-Conditioned Multi-Objective Alignment (UC-MOA), a novel framework that overcomes these limitations. Our approach leverages a diverse set of strictly increasing, non-linear utility functions to transform user-specified preferences into symbolic tokens, which are then used to condition a single LLM. This design not only mitigates numerical reasoning challenges but also substantially reduces training overhead, yielding models that achieve superior Pareto fronts and robust alignment across complex reward dimensions.
UC-MOA: Utility-Conditioned Multi-Objective Alignment for Distributional Pareto-Optimality
Zelei Chengthanks: Equal contribution. Northwestern University Evanston, USA                        Xin-Qiang Cai11footnotemark: 1 RIKEN-AIP Tokyo, Japan                        Yuting Tang The University of Tokyo Tokyo, Japan
Pushi Zhang Microsoft Research Asia Beijing, China                        Boming Yang The University of Tokyo Tokyo, Japan                        Xinyu Xing Northwestern University Evanston, USA
1 Introduction
Reinforcement Learning from Human Feedback (RLHF) has emerged as a powerful paradigm for aligning artificial agents with human values by incorporating user-provided feedback into the training process Ouyang et al. (2022); Bai et al. (2022); Stiennon et al. (2020). Despite its success, conventional RLHF approaches typically rely on scalar reward functions and struggle to capture the full spectrum of human preferences, which are inherently multi-dimensional and distributional in nature. For example, methods like RiC (Yang et al., 2024c) directly input raw reward values into the prompts. However, LLMs are known to have limited numerical reasoning capabilities, often failing to distinguish between numerical inputs (e.g., which is bigger? 9.11 versus 9.8 (Yang et al., 2024b)). This numerical insensitivity can lead to suboptimal alignment when dealing with nuanced feedback that spans multiple reward dimensions.
Concurrently, alternative strategies such as MORLHF (Li et al., 2020), Rewarded Soups (Rame et al., 2023), and MODPO (Zhou et al., 2024) attempt to address the complexity of multi-objective alignment by training multiple models to handle different reward combinations Roijers et al. (2013). While these methods provide a richer representation of human preferences, they are computationally expensive and scale poorly with the number of reward dimensions, rendering them impractical for large-scale applications. The high training cost not only limits their real-world deployment but also hinders rapid experimentation and adaptation in dynamic environments.
Figure 1: Illustration of the training processes for existing multi-objective alignment methods compared to our proposed UC-MOA. While methods such as RiC and Rewarded Soups directly inject raw reward values (which are prone to numerical errors) or require multiple models to capture diverse preferences, UC-MOA employs a single model conditioned on a symbolic token derived from a diverse set of strictly increasing utility functions.
To overcome these challenges, we propose Utility-Conditioned Multi-Objective Alignment (UC-MOA) for Distributional Pareto-Optimality. Our framework builds upon recent advances in multi-objective reinforcement learning, where the concept of Pareto-optimality is extended to distributional Pareto-optimality—a notion that prioritizes the entire return distribution over mere expectations Levy (1992); Shaked and Shanthikumar (2007). The difference of our method compared with previous methods are illustrated in Figure 1. Central to our approach is the use of a diverse set of strictly increasing, non-linear utility functions, parameterized via monotonic neural networks Amos et al. (2017); Dugas et al. (2009). These utility functions are trained on offline data with human feedback, enabling them to capture not only the expected values but also higher-order statistical properties such as variance and skewness. By converting user preferences into target rewards through a linear mapping and subsequently conditioning the LLM on a symbolic token (i.e., <max_utility_index> concatenated with the optimal utility index), UC-MOA effectively circumvents the numerical sensitivity issues that plague raw reward injection.
In our framework, the use of utility functions not only resolves the numerical reasoning challenges of previous methods but also introduces a principled mechanism to attain distributional Pareto-optimality. By ranking the outputs of these utility functions based on their percentile scores, we dynamically select the most appropriate utility perspective during training—ensuring that the resulting policy optimally balances conflicting objectives in a distributional sense. This strategy enables a single model to be conditioned on complex, multi-dimensional reward profiles, thereby dramatically reducing the training cost compared to multi-model approaches while achieving superior Pareto fronts. Our empirical evaluations demonstrate that UC-MOA robustly aligns LLM outputs with nuanced human preferences across various tasks, offering a scalable and effective solution for real-world multi-objective alignment.
2 Related Work
2.1 Large Language Model Alignment
Aligning LLMs with human preferences is a critical challenge, and RLHF has emerged as a cornerstone technique in this domain. Initially popularized by Christiano et al. (2017) using a Proximal Policy Optimization (PPO)-based framework, RLHF has played a key role in developing advanced LLMs such as ChatGPT and Claude (Bai et al., 2022; Ouyang et al., 2022). However, scaling PPO for large models introduces significant efficiency challenges (Choshen et al., 2020; Engstrom et al., 2020), motivating the development of more scalable alternatives. Recent methods like Direct Preference Optimization (DPO) (Rafailov et al., 2023) and the Reward-Ranked Fine-Tuning (RAFT) algorithm (Dong et al., 2023) specifically address alignment challenges in LLMs. DPO directly optimizes preference datasets without relying on explicit reward models, and its online iterative variants further improve exploration and robustness (Xiong et al., 2024; Liu et al., 2024). In parallel, RAFT has been successfully applied in LLMs such as LLaMA3 (Dubey et al., 2024) and Qwen2 (Yang et al., 2024a) by sampling multiple responses per prompt, ranking them via a reward model, and fine-tuning the model on the top-ranked outputs—a strategy shown to enhance reasoning capabilities (Aksitov et al., 2023; Havrilla et al., 2024).
2.2 Multi-objective Alignment
In contrast to traditional RLHF, which optimizes a single reward function, multi-objective alignment aims to balance multiple criteria simultaneously, such as harmlessness, humor, and helpfulness. This approach is crucial for real-world applications where conflicting user preferences and diverse objectives must be integrated into a model’s behavior.
Standard MORLHF (Wang et al., 2024) aggregates multiple reward signals via a linear scalarization strategy, maximizing a weighted sum of scores. However, because this method fixes the weight vector during training, it inherently supports only a single, static user preference. To overcome this limitation and better reflect the nuances of multi-objective alignment, Rewarded Soups (Rame et al., 2023) trains $N$ separate policies—each dedicated to a specific objective—and then linearly interpolates these policies. This process produces a continuous spectrum of models that can be tuned to diverse reward configurations, though it introduces scalability challenges due to the cost of training multiple policies.
Alternatively, RiC Yang et al. (2024c) adapts to different user preferences by conditioning model outputs on multiple reward signals embedded within the prompt. While this method offers flexibility, it still encounters the inherent numerical reasoning limitations of large language models, which struggle to parse and integrate raw numerical reward signals during generation.
Our proposed technique tackles the core challenges of multi-objective alignment by pre-training a diverse set of strictly increasing, non-linear utility functions. We then inject a symbolic token into the prompt that encodes the maximum utility index, enabling the model to indirectly condition its output on the reward structure without relying on precise numerical calculations. This design directly addresses the complexity of multi-objective alignment by allowing the model to seamlessly balance and integrate multiple objectives according to varied user preferences.
2.3 Distributional Reinforcement Learning.
Traditional RL focuses on estimating the expected return of a policy, but distributional reinforcement learning (DRL) extends this framework by modeling the entire distribution of returns rather than just their expectation Bellemare et al. (2017); Zhang et al. (2021). This richer representation of uncertainty has been shown to enhance both learning efficiency and policy quality in various single-objective RL tasks Dabney et al. (2018); Barth-Maron et al. (2018).
Several key algorithms have been developed in this area. Categorical DQN Bellemare et al. (2017) approximates the return distribution using a fixed set of discrete atoms, while Quantile Regression DQN Dabney et al. (2018) models it through quantile regression. Another approach, Distributional MPO Barth-Maron et al. (2018), enhances Maximum a Posteriori Policy Optimization by incorporating distributional learning, leading to improved exploration and stability.
Despite these successes, extending distributional RL to MORL remains an open challenge. The presence of multiple, often conflicting, objectives increases the complexity of reasoning about the full return distribution, as well as how preferences should be expressed over these distributions. Recent work by Cai et al. (2023) introduces Distributional Pareto-Optimal Multi-Objective Reinforcement Learning (DPMORL), which seeks to learn a set of Distributional Pareto-Optimal policies. They show that for any strictly increasing utility function, the policy maximizing expected utility corresponds to a Distributional Pareto-Optimal policy. This result suggests a principled way to incorporate distributional reasoning into MORL while preserving Pareto efficiency. Building on this theoretical foundation, our method distinguishes itself by generating a rich ensemble of non-linear utility functions using monotonic neural networks. By incorporating a symbolic selection mechanism—wherein the maximum utility index is determined and appended to the prompt—we effectively decouple the numerical complexities from the LLM’s generation process. This allows our framework to seamlessly integrate distributional considerations into the multi-objective setting while preserving the desirable properties of Pareto efficiency.
3 Proposed Technique
In this work, we propose a novel framework that integrates distributional utility modeling with RLHF. Recent work such as RiC (Yang et al., 2024d) and Rewarded Soups (Rame et al., 2023) directly appends raw reward values to the prompts. However, as argued in Section 1, LLMs exhibit limited capability in directly processing precise numerical inputs. Instead, we first train a diverse set of non-linear utility functions from an offline dataset and then introduce a symbolic token into the prompt. This token consists of the index corresponding to the maximum utility value, denoted as <max_utility_index>, concatenated with its associated utility index. During online fine-tuning, after generating a response, we determine the maximal utility index by ranking the outputs of each utility function based on their percentile scores, and then update the LLM via an online process that follows the RiC paradigm except for this novel selection mechanism. Figure 1 shows the framework of our method.
3.1 Utility Function Generation with Human Feedback
The first component of our algorithm is dedicated to constructing a rich set of candidate utility functions that will later guide the derivation of diverse optimal policies. Such diversity is crucial for addressing various distributional preferences and for adapting to different application scenarios. To achieve this, we (i) parameterize each candidate utility function using a monotonic neural network, and (ii) optimize a diversity-promoting loss that encourages discrepancies in both the output values and the local sensitivities of these functions over a normalized domain.
Monotonic Neural Network Parameterization. We denote each candidate utility function by $g_{\psi_{i}}:\mathcal{Z}\rightarrow\mathbb{R}$ for $i=1,\dots,M$, where $\mathcal{Z}\coloneqq[0,1]^{K}$ represents the normalized input space. The network $g_{\psi_{i}}$ is constructed to be monotonic (i.e., non-decreasing) by enforcing non-negativity on its weight matrices and by employing non-decreasing activation functions. This design paradigm follows the approaches proposed in convex neural networks Amos et al. (2017) and monotonic architectures such as DPMORL Cai et al. (2023), which have been shown to approximate any multivariate non-decreasing function with arbitrarily small error Dugas et al. (2009). Implementation details are provided in Appendix B.
Diversity-Promoting Objective. To ensure that the candidate functions span a wide range of utility profiles, we introduce a loss function that encourages dissimilarity both in function values and in their gradients. For a given candidate $g_{\psi_{i}}$, we first define the value discrepancy loss as
$L_{\mathrm{val}}(\psi_{i})=\min_{j\neq i}\mathbb{E}_{\mathbf{z}\sim\mathcal{U}(\mathcal{Z})}\Bigl{[}g_{\psi_{i}}(\mathbf{z})-g_{\psi_{j}}(\mathbf{z})\Bigr{]}^{2},$ (1)
where $\mathcal{U}(\mathcal{Z})$ denotes the uniform distribution over $\mathcal{Z}$. Next, to promote differences in the local behavior, we define the gradient discrepancy loss as
$\begin{split}L_{\mathrm{grad}}&(\psi_{i})=\;\min_{j\neq i}\mathbb{E}_{\begin{subarray}{c}\mathbf{z},\mathbf{z}^{\prime}\sim\mathcal{U}(\mathcal{Z})\\ \mathbf{z}\neq\mathbf{z}^{\prime}\end{subarray}}\\ &\left[\frac{g_{\psi_{i}}(\mathbf{z}^{\prime})-g_{\psi_{i}}(\mathbf{z})}{\|\mathbf{z}^{\prime}-\mathbf{z}\|}-\frac{g_{\psi_{j}}(\mathbf{z}^{\prime})-g_{\psi_{j}}(\mathbf{z})}{\|\mathbf{z}^{\prime}-\mathbf{z}\|}\right]^{2}.\end{split}$ (2)
The overall diversity loss for the candidate $g_{\psi_{i}}$ is then defined as a convex combination of these two components:
$L(\psi_{i})=\mu\,L_{\mathrm{val}}(\psi_{i})+(1-\mu)\,L_{\mathrm{grad}}(\psi_{i}),$ (3)
where the hyperparameter $\mu\in[0,1]$ balances the relative emphasis on value versus gradient discrepancies. Minimizing $L(\psi_{i})$ via gradient descent for each $i=1,\dots,M$ encourages the ensemble $\{g_{\psi_{i}}\}_{i=1}^{M}$ to cover a broad spectrum of utility profiles over $\mathcal{Z}$.
To satisfy the strict monotonicity condition required for ensuring Distributional Pareto-Optimality, we further modify each candidate by incorporating a small linear term. Specifically, we define the final utility function as
$\hat{g}_{\psi_{i}}(\mathbf{z})=g_{\psi_{i}}(\mathbf{z})+\frac{\epsilon}{K}\sum_{k=1}^{K}z_{k},$ (4)
where $\epsilon$ is a small positive constant (e.g., $\epsilon=0.01$ in our experiments). The functions $\{\hat{g}_{\psi_{i}}\}_{i=1}^{M}$ are then employed as the strictly increasing utility functions for guiding policy optimization.
3.2 Utility Index Selection during Training
To robustly determine the optimal utility index during training, we employ a percentile ranking strategy in both the offline and online phases. For every sample with reward vector $\mathbf{z}$, each candidate utility function $\hat{g}_{\psi_{i}}$ produces a utility score. We then compute the percentile rank for the $i$-th utility function as
$\text{percentile}_{i}=\frac{\text{rank of }\hat{g}_{\psi_{i}}(\mathbf{z})}{N},$ (5)
where $N$ is the total number of samples, and the rank is determined by the position of $\hat{g}_{\psi_{i}}(\mathbf{z})$ in the sorted array of scores for that function. The percentile ranking provides a normalized measure of performance for each utility function, mitigating the effects of scale differences and ensuring that selection is robust across diverse reward distributions. During offline and online training, the utility index is then selected as the one with the highest percentile rank. We present a concrete example in Appendix B.2.
3.3 Offline Training
In the offline phase, our approach establishes a robust grounding of the language model in distributional reward signals by embedding an explicit symbolic indicator of the preferred utility function directly into the training prompt. For each training sample, denoted by $(x,y)$ with $x$ representing the input prompt and $y$ the corresponding response, we first compute a multi-objective return vector $\mathbf{z}(x,y)$ and then evaluate a suite of strictly increasing candidate utility functions $\{\hat{g}_{\psi_{i}}\}_{i=1}^{M}$ on $\mathbf{z}(x,y)$. The scalar utility values $U_{i}=\hat{g}_{\psi_{i}}(\mathbf{z}(x,y))$ for $i=1,\dots,M$ are then obtained, and for each function, we calculate its percentile rank over the training distribution. The maximum utility index is determined as the index with the highest percentile:
$i^{*}=\arg\max_{i\in\{1,\dots,M\}}\left\{\text{percentile}_{i}\right\}.$ (6)
We augment the original prompt $x$ by appending a specialized token that encodes this index; that is, the modified prompt is given by
$x^{\prime}=\text{\#\#\# Prompt: }\{x\}\;\texttt{<max\_utility\_index>}+i^{*}.$
The model is then fine-tuned via a standard cross-entropy loss:
$\mathcal{L}_{\text{offline}}(\theta)=-\sum_{t}\log\pi_{\theta}(y_{t}\mid x^{\prime},y_{<t}),$
ensuring that the model learns to internalize the mapping between input contexts and their corresponding optimal distributional reward profiles. This training stage leverages supervised techniques exclusively, thereby avoiding the instability often associated with RL-based methods.
3.4 Online Fine-Tuning
To enhance the generalizability of the fine-tuned model, we incorporate additional online stages. In the online stage, the large language model generates additional candidate responses and we selectively choose those responses that exhibit superior distributional reward characteristics to fine-tune the model. Given an input prompt $x$, the model produces a response $y$ whose multi-objective return $\mathbf{z}(y)$ is computed and then evaluated by each candidate utility function, yielding scalar utility values $U_{i}=\hat{g}_{\psi_{i}}(\mathbf{z}(y))$ for $i=1,\dots,M$. For each candidate response, we compute the percentile rank of each utility score as described above, and the optimal utility index is selected as the one with the highest percentile as in Equation (5), resulting in an updated input
$x^{\prime}=\text{\#\#\# Prompt: }\{x\}\;\texttt{<max\_utility\_index>}+i^{*}.$
The response $y$ and its corresponding modified prompt $x^{\prime}$ are stored in an online training buffer for further model updates. Our online training process also incorporates a rejection sampling mechanism to filter out responses that do not exceed predetermined utility thresholds across multiple reward dimensions. The selected samples are then used to update the language model via the same reward-conditional supervised objective as in the offline training phase.
3.5 Inference Stage
At inference time, the user specifies a preference vector $\mathbf{w}=[w_{1},\ldots,w_{N}]$ that encodes the relative importance of each reward dimension $r_{1},\ldots,r_{N}$. To translate these preferences into a concrete reward profile, we apply a linear preference-to-reward mapping for each dimension:
$f_{i}(w_{i})=w_{i}\times(r_{i}^{\mathrm{max}}-r_{i}^{\mathrm{min}})+r_{i}^{\mathrm{min}},$
where $r_{i}^{\mathrm{max}}$ and $r_{i}^{\mathrm{min}}$ denote the maximum and minimum values of the $i$-th reward observed in the training dataset. This mapping yields a target reward vector $\mathbf{z}=[f_{1}(w_{1}),\dots,f_{N}(w_{N})]$ that accurately reflects the user’s preferences while normalizing the reward scales across dimensions.
Subsequently, the target reward vector $\mathbf{z}$ is input into each of the pre-trained, strictly increasing utility functions $\{\hat{g}_{\psi_{i}}\}_{i=1}^{M}$ to obtain scalar utility values $U_{i}=\hat{g}_{\psi_{i}}(\mathbf{z})$ for $i=1,\dots,M$. at inference time the reward distribution is more stable and the set of candidate responses is smaller; therefore, we directly use the argmax operation to select the index corresponding to the maximum expected utility:
$i^{*}=\arg\max_{i\in\{1,\dots,M\}}U_{i}.$ (7)
This index $i^{*}$ is encoded into the input prompt via a symbolic token—specifically, the prompt is augmented as
$x^{\prime}=\text{\#\#\# Prompt: }\{x\}\;\texttt{<max\_utility\_index>}+i^{*},$
thereby conditioning the language model to generate responses that are robustly aligned with the user’s specified multi-objective reward preferences while obviating the need for direct numerical processing.
4 Evaluation
In this section, we start with the experimental setup, including the tasks, baselines, and implementation. We then present our experiment design, followed by experiment results and analysis. We provide our code in the anonymous link https://anonymous.4open.science/r/UC-MOA-8D36.
4.1 Experiment Setup
Tasks. Following RiC (Yang et al., 2024c), we select two text generation tasks for our experiments, i.e., Helpful Assistant task (Bai et al., 2022) and the Reddit Summary task (Stiennon et al., 2020). For the Helpful Assistant task, we utilize three open-sourced reward models on Huggingface to measure three different objectives “harmless”, “helpful”, and “humor”. We focus on three multi-objective subtasks: (1) “harmless” vs “helpful”; (2) “humor” vs “helpful”; (3) “harmless” vs “humor”. Regarding the Reddit Summary task, we consider three reward models namely “summary”, “hallucinated”, and “faithful” to evaluate different perspectives of the summary quality. We focus on three multi-objective subtasks: (1) “summary” vs “faithful”; (2) “hallucinated” vs “faithful”; (3) “summary” vs “hallucinated”. We provide more details about the datasets and the reward models in Appendix A.
Baseline methods.We compare our method with two state-of-the-art multi-objective RLHF methods, Rewarded Soups, and Reward-in-Context (RiC). We also report the performance of the base model and the supervised fine-tuned (SFT) model.
Implementation. We implement the proposed method using PyTorch. We train the machine learning models on a server with 8 NVIDIA A100 80GB GPUs and 4TB memory for all methods. We select the LLaMA 2 7B-hf model as the base model for alignment following the setting of Yang et al. (2024c). We list the implementation details (e.g., network structure for the parameterized utility function, hyper-parameter choices) in Appendix B.
4.2 Experiment Design
We use the following experiments to evaluate the effectiveness and efficiency of our method, and other factors that influence the performance (e.g., alternative design choices, hyper-parameters).
Experiment I. To evaluate the effectiveness of our method, we compare the curves of the multi-dimensional average test rewards that correspond to the empirical Pareto fronts generated by different methods. A method demonstrating an outer curve indicates better performance, as it captures a wider range of trade-offs across objectives. We plot the Pareto fronts of all baseline methods across the Helpful Assistant task and the Reddit Summary task.
Experiment II. We analyze the computational efficiency of each method by recording the training time of each baseline method. We run experiments with NVIDIA A100 80GB GPUs and report the GPU hours utilized by each method.
Experiment III. To validate the necessity of the online stage, we perform an ablation study to compare the performance of pure offline training (“w/o online”), a single online iteration (“w/ online iter 1”), and two online iterations (“w/ online iter 2”).
Experiment IV. We assess the consistency of our method during the inference phase by measuring the alignment between the user-specified target objective and the reward outcomes obtained from the LLM’s generated responses. Each prompt provided to the LLM embeds a target objective requirement (e.g., in the form <max_utility_index> a, where the letter encodes the desired objective). During evaluation, the LLM is tasked with generating responses that adhere to these preset requirements. For each generated response, we compute a set of utility values using a suite of ten reward models. To account for scale differences across objectives, the computed reward values are first normalized on a per-utility basis. We then quantify consistency by comparing the expected utility and the value distribution of each utility function with the target objective extracted from the prompt.
Experiment V. To investigate how the number of utility functions impacts the performance of our method, we vary the number of utility functions from {5,10,15} and test the performance of our method with different numbers of utility functions.
Experiment VI. We investigate an alternative design choice of the utility functions, i.e., linear functions. The linear function is designed in the form of $w^{T}x$ where $||w||_{2}=1$. To construct $N$ linear utility functions, we uniformly sample $N$ points from the unit circle and set $w$ accordingly. We plot the Pareto fronts using linear utility functions and our method across all tasks.
Experiment VII. We also investigate two metrics developed specifically for assessing distributional preference, i.e., Constraint Satisfaction and Variance Objective as in previous distributional studies (Cai et al., 2023). They are designed to underscore the optimality of multivariate distributions associated with LLMs. In terms of Constraint Satisfaction, we randomly generate $M=100$ constraints. The Constraint Satisfaction metric is then computed by considering the highest probability within the policy set that satisfies each individual constraint. The Variance Objective metric, on the other hand, involves generating $M=100$ random linear weights. These weights are applied to both the expected returns and the standard deviations of returns in each dimension. This objective encourages attaining greater expected returns while simultaneously increasing variance, thereby catering to dynamic preferences. Further details about the implementation of these evaluation metrics are provided in Appendix B.4.
4.3 Experiment Results.
Performance comparison. We compare the multi-dimensional average test rewards obtained from the empirical Pareto fronts generated by different methods on two distinct tasks: the Helpful Assistant task and the Reddit Summary task. We show the results of “harmless” vs. “humor” task and “deberta” vs. “faithful” in Figure 2 while leaving the rest of the results in Appendix C.1. We observe that our method consistently produces an outer Pareto curve that dominates those of Rewarded Soups and RiC, indicating its superior ability to capture a broader spectrum of trade-offs.
Figure 2: Pareto fronts of “harmless” vs “humor” and “deberta” vs. “faithful” tasks. Each marker on the plotted curves represents the average test rewards corresponding to a particular user preference.
Computational cost. We measure the time complexity of all methods in the two-objective Helpful Assistant task and the two-objective Reddit Summary task using NVIDIA A100 80GB GPUs. We report the time cost (in the form of GPU hours) in Table 1 and Table 2. We observe that Rewarded Soups consistently consumes the most GPU hours across both tasks, while our method (w/ online iter2) achieves notably lower costs. Meanwhile, RiC requires more computing time than our method, though it remains more cost-friendly than Rewarded Soups. These findings demonstrate that our approach not only delivers strong performance (as shown in previous experiments) but also offers a clear advantage in terms of computational efficiency.
  Method GPU hours
  Rewarded Soups 69.5
RiC 20.3
Ours w/o online 8.9
Ours w/ online iter1 10.1
Ours w/ online iter2 11.3
Table 1: Comparison of GPU hours in the Helpful Assistant task, where the number of preference $M=10$, number of rewards $N=2$.
  Method GPU hours
  Rewarded Soups 15.9
RiC 5.7
Ours w/o online 2.0
Ours w/ online iter1 2.4
Ours w/ online iter2 2.8
Table 2: Comparison of GPU hours in the Reddit Summary task, where the number of preference $M=10$, number of rewards $N=2$.
Necessity of online stage. We evaluate the impact of the online stage by comparing models trained purely offline (“w/o online”) with those undergoing one (“w/ online iter 1”) and two online iterations (“w/ online iter 2”) across all tasks. We present the performance results for “harmless” vs. “humor” and “summary” vs. “faithful” tasks in Figure 3 while leaving the rest of the results in Appendix C.2. We observe that models incorporating online iterations consistently outperform the offline-only baseline. Notably, even a single online iteration yields significant improvements, with further performance gains observed when two iterations are applied, thereby underscoring the necessity of the online stage in our training framework. The online stage is important because it enables the model to generalize over unseen samples in the offline dataset, and ultimately enhancing its ability to generate appropriate responses corresponding to the required utility function index.
Figure 3: Ablation study of the online stage on “harmless” vs. “humor” task and “summary” vs. “faithful” task.
Consistency. We conduct a case study on the “harmless” vs. “humor” task, appending the requirement “<max_utility_idx> i” to each prompt to encourage the model to follow the first preference (i.e., maximizing 9th utility function). We then use reward models to evaluate the generated responses. We provide a detailed analysis in Appendix C.3 to present the distribution of normalized utility values for each utility function over our evaluation data. Notably, the 9th utility function achieves the highest mean normalized value, indicating that the fine-tuned model aligns its responses with the specified preference requirement. This demonstrates the model’s ability to selectively prioritize and adhere to user-designated objectives.
Number of utility functions. We vary the number of utility functions over the set 5, 10, 15 and compare the resulting empirical Pareto fronts in Figure 4. We leave the results of the remaining four tasks in Appendix C.4. With 5 utility functions, the generated Pareto front is slightly inferior to those produced with 10 and 15 utility functions. Specifically, the front with 5 utility functions captures a narrower range of optimal trade-offs across the objectives. However, when the number of utility functions is increased to 10 or 15, the resulting Pareto fronts exhibit comparable performance. Increasing the number of utility functions beyond 10 does not yield substantial gains. These results demonstrate that our method is insensitive to the precise number of utility functions, provided that a sufficient number (around 10) is used to effectively cover the preference space. In practical terms, employing 10 utility functions strikes a good balance between performance and computational efficiency, as further increases to 15 do not produce notable improvements.
Figure 4: Ablation study of different numbers of utility functions on “humor” vs. “helpful” and “deberta” vs. “faithful” tasks.
Variants of utility functions. We explored the impact of using linear utility functions, designed as $w^{T}x$ where $||w||_{2}=1$, and compared the performance with our original approach across all tasks. The generated Pareto fronts, as shown in Figure 5 and Appendix C.5, illustrate that our method continues to deliver strong results even with linear utility functions. However, there is still a performance gap between the results achieved with distributional Pareto-optimal utility functions and those obtained using linear functions. While linear functions still produce reasonably good trade-offs, they fail to capture the full range of optimal solutions attainable by our method, which leverages distributional utility representations. This highlights the advantage of our approach in handling more complex utility functions, demonstrating its superior flexibility in navigating multi-objective optimization.
Figure 5: Ablation study of using linear utility functions.
Constraint satisfaction and variance objective. In our case study on the “harmless” vs. “humor” task, we use 2,000 inference samples and compute both metrics for our method, Rewarded Soup, and RiC. As summarized in Table 3, our method achieves the highest overall constraint satisfaction ratio and exhibits the largest variance objective value, outperforming both Rewarded Soup and RiC. This indicates that our approach offers a more comprehensive distributional coverage, effectively handling diverse user preferences.
  Method Constraint Satisfaction Variance
RiC 0.5637 0.0662
Rewarded Soups 0.6744 0.1631
Ours 0.7184 0.5160
Table 3: Comparison of Methods for Constraint satisfaction and variance objective.
5 Conclusion
We propose Utility-Conditioned Multi-Objective Alignment (UC-MOA), a novel multi-objective LLM alignment framework. By transforming user-specified preference into symbolic tokens through a diverse ensemble of strictly increasing, non-linear utility functions, UC-MOA circumvents the numerical sensitivity issues seen in methods like RiC while avoiding the high computational costs of approaches such as MORLHF, Rewarded Soups, and MODPO. Our experiments demonstrate that UC-MOA achieves superior Pareto fronts across complex reward dimensions, evidencing robust and scalable alignment with human values.
Limitations
Representation of Human Preferences. Although UC-MOA effectively captures the distributional properties of human feedback, its performance heavily depends on the quality and diversity of the underlying preference data. Real-world human feedback is often noisy or biased, and addressing such discrepancies requires robust preprocessing and annotation methods. Additionally, certain complex preferences might not be adequately represented by the utility functions, potentially limiting the model’s applicability in diverse contexts, such as safety-critical domains like healthcare and autonomous systems.
Computational and Scaling Challenges. The proposed framework improves efficiency compared to existing methods, but it still requires substantial computational resources when scaling to tasks with numerous objectives or high-dimensional reward spaces. The reliance on large-scale GPUs and the requirement for iterative online updates may limit its adoption in settings with constrained computational budgets. Further optimization of the model architecture and training process could help mitigate these challenges, making the framework more accessible and enabling faster experimentation.
Generalizability Across Tasks. While UC-MOA has demonstrated superior alignment performance in text generation and summary tasks, its applicability to other domains, such as multimodal tasks, remains unexplored. Extending the framework to these areas might introduce additional challenges, including the need to handle dynamic environments and integrate feedback in near real-time. Adapting and validating this approach across a broader range of use cases to evaluate its generalizability and robustness is a potential direction for future research.
Ethical Considerations
While multi-objective optimization and reward modeling can enhance the alignment of large language models, our method must be applied responsibly to avoid harmful outcomes. The potential biases or harmfulness in training data or reward models may mislead the final LLM’s behavior. Careful curation of training and evaluation datasets, combined with auditing of model outputs, is essential to mitigate these risks.
Appendix A Details of Datasets
In this section, we introduce the HH-RLHF dataset and Reddit Summary dataset in more detail. We provide the links to the HH-RLHF dataset and the three reward models we used for the experiments in Table 4.
HH-RLHF Dataset
Dataset Anthropic/hh-rlhf
Harmless Reward gpt2-large-harmless-reward_model
Helpful Reward gpt2-large-helpful-reward_model
Humor Reward humor-no-humor
Reddit Summary Dataset
Dataset openai/summarize_from_feedback
Summary Reward gpt2_reward_summarization
Deberta Reward reward-model-deberta-v3-large-v2
Faithful Reward bart-faithful-summary-detector
Table 4: Datasets and Reward Models.
HH-RLHF Dataset
The HH-RLHF dataset, as described in Bai et al. (2022), was collected to support research on aligning language models using human preferences. The dataset comprises approximately 160,000 prompts each paired with two responses generated by a language model. For every prompt, one response is labeled as “chosen” (preferred by human annotators) and the other as “rejected”. Human raters evaluated paired responses based on criteria for helpfulness and harmlessness. For helpfulness, annotators selected the response that best addressed the given task, whereas for harmlessness (red-teaming) they identified the response as more likely to produce harmful content. This dual-annotation design allows for training reward models that capture subtle aspects of human judgment. Note that we follow Yang et al. (2024d) and use “chosen” data to construct our offline dataset.
Reddit Summary Dataset
The Reddit Summary dataset, introduced in Stiennon et al. (2020), is intended for research in abstractive summarization with human feedback. The dataset consists of approximately 14,900 Reddit posts paired with corresponding summaries. Each entry comprises a post and its summary, where the summary is either the original TL;DR provided by the poster or a human-annotated summary aimed at capturing the post’s main ideas.
Appendix B Implementation Details
B.1 Utility Functions
Non-Decreasing Neural Network.
We used a Non-Decreasing Neural Network to parameterize the utility function. The network architecture consisted of three fully connected layers, with our designed non-decreasing activation functions applied to the hidden layers:
$\begin{split}f(x)=\text{concat}[&max(x,-0.5),min(x,0.5),\\ &clip(x,-0.5,0.5)].\end{split}$ (8)
This design of the activation function is an extension of the ReLU activation function, which empirically makes the learned utility function more diverse, covering more types of decision boundaries. After each gradient update during training, we clip all the trainable weights in the network to be non-negative to ensure that the resulting utility function was non-decreasing. The network’s input dimension corresponded to the number of objectives in the environment, and the output dimension was a scalar representing the utility value.
Return Normalization.
We use a normalization technique by linearly mapping the return into scale $[0,1]^{K}$ without modifying the optimal policies. Specifically, we keep track of the minimum and maximum multivariate returns that the policies have currently achieved, denoted by $\mathbf{z}_{min}$ and $\mathbf{z}_{max}$. We let $\mathbf{z}_{mid}=(\mathbf{z}_{min}+\mathbf{z}_{max})/2$, and $d=\max_{i}[\mathbf{z}_{max,i}-\mathbf{z}_{min,i}]$. Before any multivariate return $z$ is fed into the utility function, we normalize it by
$z_{norm}=\frac{1}{d}(z_{norm}-z_{mid})+\frac{1}{2}$
This maps $z$ linearly to $z_{norm}$, where $z_{norm}\in[0,1]^{K}$, without affecting the scale of each dimension of returns. In the implementation, the return $\mathbf{z}_{t}$ is normalized before feeding into the utility function to compute the scalar reward.
B.2 Prompt Labeling
We provide an example of our prompt labeling technique in the “harmless” vs. “helpful” task. As illustrated in Figure 6, the response is evaluated in two dimensions “harmless” and “helpful”. The two-dimensional scores are further feed into ten utility functions and obtain a set of utility scores. We then convert these scores into percentile rankings. Finally, we label the prompt with the utility function whose percentile ranking is highest. In the example, the highest percentile ranking belongs to utility index 2 (67.67%), so we append the prompt with <max_utility_idx> c. This procedure ensures that each prompt-response pair is labeled with the most prominent utility dimension, i.e., the one where its performance stands out most relative to other examples, thereby facilitating more focused downstream alignment.
Figure 6: An example of our prompt labeling technique in “harmless” vs. “helpful” dataset.
B.3 Hyperparameter Choices
Table 5 summarizes the hyper-parameter setting for our experiments.
Training Phase
Offline Learning Rate $1.414\times 10^{-4}$
Online Learning Rate $10^{-5}$
Batch Size assistant: $16$
summary: $2$
Online Generation kwargs max_new_tokens (assistant): $128$
max_new_tokens (summary): $48$
num_of_samples: $1000$
min_length: $-1$
top_k: 00
top_p: $0.9$
temperature: $0.9$
do_sample: True
begin_suppress_tokens: [tokenizer.eos_token_id]
Evaluation Phase
Generation kwargs max_new_tokens (assistant): $128$
max_new_tokens (summary): $48$
min_length: $-1$
top_k: $0.0$
top_p: $0.9$
do_sample: True
Table 5: Hyper-parameter settings in our experiments.
B.4 Distributional Metrics
We use two novel metrics as in Cai et al. (2023) for assessing distributional preference: Constraint Satisfaction and Variance Objective.
For Constraint Satisfaction, we define $u(p,\pi)$ to be the probability of return random variance $z\sim\mathbf{Z}(\pi)$ that satisfies the constraint $p$, representing how a policy satisfies the given constraint. We randomly generate each constraint defined by $n$ linear constraints represented by $\{\mathbf{z}|\forall i,\mathbf{w}_{i}^{T}\mathbf{z}\geq c_{i}\}$, where $w_{i}$ is uniformly sampled from the surface $\mathbf{w}_{i}\succeq 0,\mathbf{1}^{T}\mathbf{w}_{i}=1$, and $c_{i}$ is sampled from the distribution of $\mathbf{U}(\min_{\mathbf{z}\in\mathbf{Z}}\mathbf{w}_{i}^{T}\mathbf{z},\max_{\mathbf{z}\in\mathbf{Z}}\mathbf{w}_{i}^{T}\mathbf{z})$, where $\mathbf{Z}$ is the set of all return samples of the final policies learned by DPMORL.
For Variance Objective, we define $u(p,\pi)$ to be a weighted sum of expected return $\mathbb{E}_{\mathbf{z}\sim\mathbf{\mu}(\pi)}\mathbf{z}$ and standard deviation of return $\sqrt{\text{Var}_{\mathbf{z}\sim\mathbf{\mu}(\pi)}[\mathbf{z}]}$. We randomly generate $\mathbf{w}_{i}\in\mathbb{R}^{2K}$, where $\mathbf{w}_{i}\succeq 0$ and $\mathbf{1}^{T}\mathbf{w}_{i}=1$. We let $\mathbf{w}_{i}=[\mathbf{w}_{i}^{1},\mathbf{w}_{i}^{2}]$, where $\mathbf{w}_{i}^{1},\mathbf{w}_{i}^{2}\in\mathbb{R}^{K}$, and use $\mathbf{w}_{i}^{1}$ and -$\mathbf{w}_{i}^{2}$ as the weight for expected return and standard derivation of return respectively.
For both of the metrics, we define $u(p,\pi)$ to be the satisfaction score of policy $\pi$ under preference $p$. We generate $M=100$ number of preference $p_{1},p_{2},\cdots,p_{M}$. For both of the Constraint Satisfaction and Variance Objective metric, the mean score is defined as
$\frac{1}{M}\sum_{i=1}^{M}\max_{j\in\{1,2,\cdots,N\}}u(p_{i},\pi_{j})$
The mean score in the above equation measures the best performance of satisfying the preference of the user, averaged in $M=100$ preferences. A higher mean score represents a better ability to meet different preferences.
Appendix C Additional Experiment Results
C.1 Experiment I
In addition to the main results presented in the paper, Figure 7 shows the Pareto fronts for the remaining four tasks in our experiments. We compare five methods: Llama 2 base, SFT, Rewarded Soups, RiC, and our proposed approach (Ours). Each subplot corresponds to a different two-objective trade-off, where the $x$- and $y$-axes represent the average test rewards for two distinct objectives. A method demonstrating an outer Pareto curve indicates a more favorable performance, as it captures a broader spectrum of trade-offs between competing objectives. Across all four subplots in Figure 7, Our method (purple line) achieves the most outward Pareto front, reflecting superior coverage of the objective space compared to the other baselines. RiC (orange line) also performs strongly but is dominated by Ours in several regions of the objective space. Meanwhile, Rewarded Soups (red line), SFT (green crosses), and Llama 2 base (blue triangles) display varying degrees of suboptimality in balancing the two objectives simultaneously. Overall, these results reinforce our main findings, showing that our method provides a robust and effective means of navigating multi-objective trade-offs across a range of different tasks.
Figure 7: Pareto fronts of the remaining four tasks.
C.2 Experiment III
We investigate the impact of the online training stage by comparing three variants of our method: (1) pure offline training (w/o online), (2) one online iteration (w/ online iter 1), and (3) two online iterations (w/ online iter 2). As shown in Figure 8, each subplot corresponds to the Pareto fronts for a different task. Across all four tasks, w/o online (blue line) consistently produces the weakest Pareto front, indicating suboptimal performance when balancing multiple objectives. Introducing a single online iteration (w/ online iter 1, green line) noticeably expands the Pareto front, suggesting that even a brief online training phase can significantly improve multi-reward optimization. Finally, adding a second online iteration (w/ online iter 2, red line) further refines these trade-offs, achieving the most outward (and thus superior) Pareto front among the three variants. These results confirm the necessity of the online stage, demonstrating that multiple online iterations can better align the model with the target objectives.
Figure 8: Ablation study of the online stage on the remaining four tasks.
C.3 Experiment IV
We examine how closely our model’s generated responses align with a user-specified target objective, in this case, the 9th utility function (index 8). Figure 9 shows the distribution of normalized utility values across all ten reward models for the “harmless” vs. “humor” task, and Table 6 reports the corresponding mean normalized values. Notably, the 9th utility function (index 8) attains the highest average score of 0.68, demonstrating that the LLM effectively prioritizes the user-indicated objective. In other words, when prompts include the requirement <max_utility_index> i specifying the 9th utility, the model consistently generates responses that yield higher scores along that dimension than along the others. These findings confirm that our fine-tuned model can reliably follow the user-designated target objective during inference.
Figure 9: Utility value distribution for the “harmless” vs. “humor” task. The required utility index is 8 (i.e., the 9th utility function).
Utility index 0 1 2 3 4 5 6 7 8 9
Mean 0.57 0.57 0.56 0.56 0.57 0.57 0.49 0.55 0.68 0.58
Table 6: Mean normalized value for each utility function.
C.4 Experiment V
We investigate how varying the number of utility functions (5, 10, 15) affects our method’s ability to balance multiple objectives. Figure 10 shows the resulting Pareto fronts across four different tasks, where an outward-shifted curve indicates better overall performance. We observe that using only 5 utility functions (blue line) yields a slightly weaker Pareto front compared to the other two settings. In contrast, the results for 10 (green line) and 15 (red line) utility functions are nearly identical, suggesting that increasing beyond 10 utility functions does not lead to a substantial performance gain. Consequently, our method demonstrates robustness to the exact number of utility functions, so long as it is sufficiently large to capture the relevant trade-offs among the objectives.
Figure 10: Ablation study of different numbers of utility functions on remaining four tasks.
C.5 Experiment VI
Figure 11: Abalation study of using linear utility functions.
We compare the performance of our method when using linear utility functions (Ours (linear)) versus our original, distribution-based approach (Ours (param)) in Figure 11. In each subplot, the curves depict how effectively each method covers the trade-offs between two objectives. Although the linear approach maintains a reasonable Pareto front and thus remains effective for balancing multiple rewards, it consistently trails the distribution-based version in all tasks. This shortfall indicates that purely linear utility functions, while simpler to construct, capture a narrower range of potential trade-offs and consequently yield a less comprehensive frontier. In contrast, our distribution-based design is more flexible, enabling better alignment with diverse user preferences and resulting in a noticeably more outward Pareto front.

### v2

@@
 Abstract
-Reinforcement Learning from Human Feedback (RLHF) has become a cornerstone for aligning large language models (LLMs) with human values.
-However, existing approaches struggle to capture the multi-dimensional, distributional nuances of human preferences.
-Methods such as RiC that directly inject raw reward values into prompts face significant numerical sensitivity issues—for instance, LLMs may fail to distinguish between 9.11 and 9.8—while alternatives like MORLHF, Rewarded Soups, and MODPO incur high computational costs by training multiple models.
-In this work, we introduce Utility-Conditioned Multi-Objective Alignment (UC-MOA), a novel framework that overcomes these limitations.
-Our approach leverages a diverse set of strictly increasing, non-linear utility functions to transform user-specified preferences into symbolic tokens, which are then used to condition a single LLM.
-This design not only mitigates numerical reasoning challenges but also substantially reduces training overhead, yielding models that achieve superior Pareto fronts and robust alignment across complex reward dimensions.
-UC-MOA: Utility-Conditioned Multi-Objective Alignment for Distributional Pareto-Optimality
-Zelei Chengthanks: Equal contribution.
-Northwestern University Evanston, USA                        Xin-Qiang Cai11footnotemark: 1 RIKEN-AIP Tokyo, Japan                        Yuting Tang The University of Tokyo Tokyo, Japan
-Pushi Zhang Microsoft Research Asia Beijing, China                        Boming Yang The University of Tokyo Tokyo, Japan                        Xinyu Xing Northwestern University Evanston, USA
+Aligning Large Language Models (LLMs) to individual user needs is paramount for effective personalization.
+However, capturing the diverse and often nuanced preferences of different users presents a significant challenge.
+Existing approaches may not adequately represent complex preference trade-offs, can be computationally expensive, or rely on the LLM’s limited numerical reasoning capabilities for personalization.
+We introduce Utility-Conditioned Multi-Objective Alignment (UC-MOA), a novel framework for personalized LLM alignment.
+UC-MOA develops a diverse set of utility functions, each representing a distinct preference structure, to cater to a wide spectrum of user requirements.
+These utility functions are constructed with a theoretical guarantee to ensure the resulting set of preference-aligned policies can achieve distributional Pareto-optimality.
+Instead of conditioning on raw numerical values, UC-MOA employs a “utility-conditioned” method, using a symbolic token that represents an optimal utility function index.
+This allows the LLM to learn an implicit mapping between the token and the desired personalized response characteristics through fine-tuning, thereby enabling effective personalization.
+We conduct a user study demonstrating UC-MOA’s superior ability to satisfy diverse user preferences compared to baseline methods, and perform quantitative experiments showing that UC-MOA achieves superior Pareto fronts and maintains high computational efficiency, offering a scalable and robust solution for multi-objective personalized LLM alignment.
 1 Introduction
-Reinforcement Learning from Human Feedback (RLHF) has emerged as a powerful paradigm for aligning artificial agents with human values by incorporating user-provided feedback into the training process Ouyang et al.
-(2022); Bai et al.
-(2022); Stiennon et al.
-(2020).
-Despite its success, conventional RLHF approaches typically rely on scalar reward functions and struggle to capture the full spectrum of human preferences, which are inherently multi-dimensional and distributional in nature.
-For example, methods like RiC (Yang et al., 2024c) directly input raw reward values into the prompts.
-However, LLMs are known to have limited numerical reasoning capabilities, often failing to distinguish between numerical inputs (e.g., which is bigger?
-9.11 versus 9.8 (Yang et al., 2024b)).
-This numerical insensitivity can lead to suboptimal alignment when dealing with nuanced feedback that spans multiple reward dimensions.
-Concurrently, alternative strategies such as MORLHF (Li et al., 2020), Rewarded Soups (Rame et al., 2023), and MODPO (Zhou et al., 2024) attempt to address the complexity of multi-objective alignment by training multiple models to handle different reward combinations Roijers et al.
-(2013).
-While these methods provide a richer representation of human preferences, they are computationally expensive and scale poorly with the number of reward dimensions, rendering them impractical for large-scale applications.
-The high training cost not only limits their real-world deployment but also hinders rapid experimentation and adaptation in dynamic environments.
-Figure 1: Illustration of the training processes for existing multi-objective alignment methods compared to our proposed UC-MOA.
-While methods such as RiC and Rewarded Soups directly inject raw reward values (which are prone to numerical errors) or require multiple models to capture diverse preferences, UC-MOA employs a single model conditioned on a symbolic token derived from a diverse set of strictly increasing utility functions.
-To overcome these challenges, we propose Utility-Conditioned Multi-Objective Alignment (UC-MOA) for Distributional Pareto-Optimality.
-Our framework builds upon recent advances in multi-objective reinforcement learning, where the concept of Pareto-optimality is extended to distributional Pareto-optimality—a notion that prioritizes the entire return distribution over mere expectations Levy (1992); Shaked and Shanthikumar (2007).
-The difference of our method compared with previous methods are illustrated in Figure 1.
-Central to our approach is the use of a diverse set of strictly increasing, non-linear utility functions, parameterized via monotonic neural networks Amos et al.
-(2017); Dugas et al.
-(2009).
-These utility functions are trained on offline data with human feedback, enabling them to capture not only the expected values but also higher-order statistical properties such as variance and skewness.
-By converting user preferences into target rewards through a linear mapping and subsequently conditioning the LLM on a symbolic token (i.e., <max_utility_index> concatenated with the optimal utility index), UC-MOA effectively circumvents the numerical sensitivity issues that plague raw reward injection.
-In our framework, the use of utility functions not only resolves the numerical reasoning challenges of previous methods but also introduces a principled mechanism to attain distributional Pareto-optimality.
-By ranking the outputs of these utility functions based on their percentile scores, we dynamically select the most appropriate utility perspective during training—ensuring that the resulting policy optimally balances conflicting objectives in a distributional sense.
-This strategy enables a single model to be conditioned on complex, multi-dimensional reward profiles, thereby dramatically reducing the training cost compared to multi-model approaches while achieving superior Pareto fronts.
-Our empirical evaluations demonstrate that UC-MOA robustly aligns LLM outputs with nuanced human preferences across various tasks, offering a scalable and effective solution for real-world multi-objective alignment.
+Large Language Models (LLMs) have demonstrated impressive capabilities across a wide range of natural language tasks and increasingly serve as powerful tools for real-world applications [1, 2, 3].
+Developing personalized LLMs are essential for business monetization, as it enables tailored user experiences that enhance engagement, retention, and revenue [4, 5, 6].
+However, achieving effective personalization is non-trivial, as it requires addressing diverse and often conflicting user preferences, which cannot be adequately represented by a single scalar reward [7, 8].
+Conventional fine-tuning approaches, such as Proximal Policy Optimization (PPO) [9] and Direct Preference Optimization (DPO) [10] optimize a single or aggregated reward signal, which fails to capture the multi-objective nature of real-world user requests.
+Consequently, there is an urgent need for multi-objective alignment methods that can balance multiple goals to satisfy diverse individual preferences.
+Existing methods for multi-objective alignment in LLMs face challenges when applied to fine-grained personalization.
+Traditional Multi-Objective Reinforcement Learning from Human Feedback (MORLHF) uses linear scalarization of reward signals [11].
+This approach struggles to capture complex, non-linear preference structures unique to individual users and typically requires training separate models for different preference weightings, making it computationally prohibitive for wide-scale personalization.
+More flexible techniques like Rewarded Soups [12] interpolate between multiple fine-tuned models, but still incur the high cost of training and storing multiple LLMs.
+A recent alternative, Reward-in-Context (RiC) [13] seeks to achieve personalization within a single model by embedding numerical reward targets directly into the prompt, thereby guiding the LLM toward a user-specified point in the reward space.
+However, a significant drawback of RiC is their reliance on a continuous reward space.
+To achieve satisfactory personalization performance, the LLM must possess robust numerical reasoning capabilities [14].
+Specifically, it needs to accurately judge which training examples or potential generations are quantitatively closer to the user’s multi-dimensional preference vector.
+This dependence on precise numerical interpretation can be a bottleneck, as LLMs often struggle with such fine-grained quantitative discernment [15], potentially leading to inconsistent or suboptimal personalization outcomes.
+To address these limitations and advance personalized LLM alignment, we propose Utility-Conditioned Multi-Objective Alignment (UC-MOA).
+At the core of UC-MOA is the development of a diverse set of non-linear utility functions, each intended to capture a distinct user preference structure across multi-dimensional reward outcomes.
+Importantly, this diverse set of utility functions is constructed with theoretical guarantees ensuring that the policies aligned with these utilities can achieve distributional Pareto-optimality [16].
+This means the framework can identify policies that offer optimal trade-offs across objectives, considering the entire distribution of outcomes, without being dominated by other achievable policies.
+Instead of injecting numerical reward targets, which can tax the LLM’s numerical reasoning, we employ a symbolic conditioning mechanism.
+For each response, we determine which utility function in our diverse set it best satisfies.
+We then associate this response with a discrete symbolic token, <max_utility_index> $i^{*}$, where $i^{*}$ is the index of that optimal utility function.
+Through fine-tuning, the LLM learns an implicit mapping between this symbolic token and the nuanced characteristics of responses that align with the corresponding underlying utility function.
+This strategy enables fine-grained personalization according to diverse preference profiles while sidestepping the LLM’s numerical reasoning limitations.
+To summarize, our contributions are:
+  • UC-MOA introduces a novel method for personalized LLM alignment by training and leveraging a diverse set of utility functions to simulate varied user preferences.
+These utility functions are designed to ensure the resulting set of aligned policies can achieve distributional Pareto-optimality.
+  • We propose a symbolic utility-conditioned mechanism where the LLM learns a mapping from utility indices to desired response styles, enabling robust personalization while avoiding the pitfalls of direct numerical reward injection.
+  • Our comprehensive experiments demonstrate the superiority of UC-MOA in satisfying user preference, achieving competitive Pareto fronts for diverse preference profiles, and maintaining high computational efficiency, offering a scalable and effective solution for personalizing LLMs to complex, multi-faceted user preferences.
 2 Related Work
-2.1 Large Language Model Alignment
-Aligning LLMs with human preferences is a critical challenge, and RLHF has emerged as a cornerstone technique in this domain.
-Initially popularized by Christiano et al.
-(2017) using a Proximal Policy Optimization (PPO)-based framework, RLHF has played a key role in developing advanced LLMs such as ChatGPT and Claude (Bai et al., 2022; Ouyang et al., 2022).
-However, scaling PPO for large models introduces significant efficiency challenges (Choshen et al., 2020; Engstrom et al., 2020), motivating the development of more scalable alternatives.
-Recent methods like Direct Preference Optimization (DPO) (Rafailov et al., 2023) and the Reward-Ranked Fine-Tuning (RAFT) algorithm (Dong et al., 2023) specifically address alignment challenges in LLMs.
-DPO directly optimizes preference datasets without relying on explicit reward models, and its online iterative variants further improve exploration and robustness (Xiong et al., 2024; Liu et al., 2024).
-In parallel, RAFT has been successfully applied in LLMs such as LLaMA3 (Dubey et al., 2024) and Qwen2 (Yang et al., 2024a) by sampling multiple responses per prompt, ranking them via a reward model, and fine-tuning the model on the top-ranked outputs—a strategy shown to enhance reasoning capabilities (Aksitov et al., 2023; Havrilla et al., 2024).
-2.2 Multi-objective Alignment
-In contrast to traditional RLHF, which optimizes a single reward function, multi-objective alignment aims to balance multiple criteria simultaneously, such as harmlessness, humor, and helpfulness.
-This approach is crucial for real-world applications where conflicting user preferences and diverse objectives must be integrated into a model’s behavior.
-Standard MORLHF (Wang et al., 2024) aggregates multiple reward signals via a linear scalarization strategy, maximizing a weighted sum of scores.
-However, because this method fixes the weight vector during training, it inherently supports only a single, static user preference.
-To overcome this limitation and better reflect the nuances of multi-objective alignment, Rewarded Soups (Rame et al., 2023) trains $N$ separate policies—each dedicated to a specific objective—and then linearly interpolates these policies.
-This process produces a continuous spectrum of models that can be tuned to diverse reward configurations, though it introduces scalability challenges due to the cost of training multiple policies.
-Alternatively, RiC Yang et al.
-(2024c) adapts to different user preferences by conditioning model outputs on multiple reward signals embedded within the prompt.
-While this method offers flexibility, it still encounters the inherent numerical reasoning limitations of large language models, which struggle to parse and integrate raw numerical reward signals during generation.
-Our proposed technique tackles the core challenges of multi-objective alignment by pre-training a diverse set of strictly increasing, non-linear utility functions.
-We then inject a symbolic token into the prompt that encodes the maximum utility index, enabling the model to indirectly condition its output on the reward structure without relying on precise numerical calculations.
-This design directly addresses the complexity of multi-objective alignment by allowing the model to seamlessly balance and integrate multiple objectives according to varied user preferences.
-2.3 Distributional Reinforcement Learning.
-Traditional RL focuses on estimating the expected return of a policy, but distributional reinforcement learning (DRL) extends this framework by modeling the entire distribution of returns rather than just their expectation Bellemare et al.
-(2017); Zhang et al.
-(2021).
-This richer representation of uncertainty has been shown to enhance both learning efficiency and policy quality in various single-objective RL tasks Dabney et al.
-(2018); Barth-Maron et al.
-(2018).
-Several key algorithms have been developed in this area.
-Categorical DQN Bellemare et al.
-(2017) approximates the return distribution using a fixed set of discrete atoms, while Quantile Regression DQN Dabney et al.
-(2018) models it through quantile regression.
-Another approach, Distributional MPO Barth-Maron et al.
-(2018), enhances Maximum a Posteriori Policy Optimization by incorporating distributional learning, leading to improved exploration and stability.
-Despite these successes, extending distributional RL to MORL remains an open challenge.
-The presence of multiple, often conflicting, objectives increases the complexity of reasoning about the full return distribution, as well as how preferences should be expressed over these distributions.
-Recent work by Cai et al.
-(2023) introduces Distributional Pareto-Optimal Multi-Objective Reinforcement Learning (DPMORL), which seeks to learn a set of Distributional Pareto-Optimal policies.
-They show that for any strictly increasing utility function, the policy maximizing expected utility corresponds to a Distributional Pareto-Optimal policy.
-This result suggests a principled way to incorporate distributional reasoning into MORL while preserving Pareto efficiency.
-Building on this theoretical foundation, our method distinguishes itself by generating a rich ensemble of non-linear utility functions using monotonic neural networks.
-By incorporating a symbolic selection mechanism—wherein the maximum utility index is determined and appended to the prompt—we effectively decouple the numerical complexities from the LLM’s generation process.
-This allows our framework to seamlessly integrate distributional considerations into the multi-objective setting while preserving the desirable properties of Pareto efficiency.
+2.1 Multi-Objective Personalization of LLMs
+Traditional LLM alignment, using methods such as PPO [17, 1, 2] and DPO [10], primarily optimizes for a single reward signal.
+This is ill-suited to the realities of personalized assistants, where users care about several, often conflicting, attributes simultaneously [18].
+To address this, MORLHF [11] extends PPO for multiple objectives, often by treating user preferences as different weighting schemes.
+However, its common reliance on linear scalarization struggles with complex, non-linear preference trade-offs, and scaling to many unique user profiles by training separate models is impractical.
+Rewarded Soups [12] offer parameter-merging strategies but still face scalability challenges for fine-grained, diverse personalization.
+More direct conditioning approaches aim to achieve personalization within a single model.
+RiC [13] attempts to condition the LLM directly on reward preferences by embedding target reward values into the input prompt.
+For instance, the model may be instructed to generate responses with specified numerical scores for attributes such as helpfulness and harmlessness.
+However, the effectiveness of RiC relies heavily on the LLM’s capacity for numerical reasoning, particularly in accurately interpreting continuous input signals, which can present a significant limitation for reliable personalization.
+Our UC-MOA framework offers a distinct approach to personalized multi-objective alignment.
+It begins by learning a library of strictly non-decreasing, non-linear utility functions, each capturing a unique preference structure.
+The monotonicity constraint ensures that increases in any individual reward dimension do not decrease the overall utility, thereby preserving consistency with rational user behavior.
+Rather than conditioning the model on numerical reward targets, UC-MOA uses symbolic tokens representing the selected utility functions to guide a single fine-tuned model.
+This circumvents the LLM’s numerical reasoning limitations while still enabling fine-grained personalization across diverse multi-objective preferences.
+2.2 Distributional Pareto-Optimality
+Effective multi-objective personalization requires more than optimizing expected reward values, as users often have preferences concerning the distribution of outcomes (e.g., risk, consistency) across objectives [19, 2].
+Distributional Pareto-Optimal Multi-Objective RL (DPMORL) [16] provides a formal grounding for distributional Pareto-optimality.
+DPMORL defines Pareto optimality over return distributions, where a policy is optimal if no other policy stochastically dominates it.
+Critically, DPMORL shows that policies maximizing any strictly increasing utility function over the multi-objective return distribution are part of the Distributional Pareto set.
+This links diverse utility functions to the exploration of distributionally optimal trade-offs.
+UC-MOA leverages this by constructing a diverse set of utility functions designed to span a comprehensive range of preference structures over distributional outcomes.
+By conditioning the LLM on symbolic indices tied to these utility functions, UC-MOA implicitly guides the model towards policies that are distributionally Pareto-optimal.
+This ensures that the achieved personalization is not only diverse but also adheres to a rigorous standard of optimality regarding the full spectrum of possible outcome distributions, without requiring the LLM to perform explicit distributional calculations.
 3 Proposed Technique
-In this work, we propose a novel framework that integrates distributional utility modeling with RLHF.
-Recent work such as RiC (Yang et al., 2024d) and Rewarded Soups (Rame et al., 2023) directly appends raw reward values to the prompts.
-However, as argued in Section 1, LLMs exhibit limited capability in directly processing precise numerical inputs.
-Instead, we first train a diverse set of non-linear utility functions from an offline dataset and then introduce a symbolic token into the prompt.
-This token consists of the index corresponding to the maximum utility value, denoted as <max_utility_index>, concatenated with its associated utility index.
-During online fine-tuning, after generating a response, we determine the maximal utility index by ranking the outputs of each utility function based on their percentile scores, and then update the LLM via an online process that follows the RiC paradigm except for this novel selection mechanism.
-Figure 1 shows the framework of our method.
-3.1 Utility Function Generation with Human Feedback
-The first component of our algorithm is dedicated to constructing a rich set of candidate utility functions that will later guide the derivation of diverse optimal policies.
-Such diversity is crucial for addressing various distributional preferences and for adapting to different application scenarios.
-To achieve this, we (i) parameterize each candidate utility function using a monotonic neural network, and (ii) optimize a diversity-promoting loss that encourages discrepancies in both the output values and the local sensitivities of these functions over a normalized domain.
+Effectively personalizing LLMs to diverse user preferences across multiple objectives is a significant challenge.
+While recent methods such as RiC [20] attempt personalization by appending raw numerical reward values directly to prompts, LLMs often struggle with precise numerical inputs, limiting their ability to reliably adapt to nuanced user needs, as discussed in Section 1.
+Our proposed framework, UC-MOA, circumvents this limitation to achieve robust multi-objective personalization.
+UC-MOA integrates distributional utility modeling with RLHF.
+The core idea is to first learn a diverse library of non-linear, strictly increasing utility functions, each representing a distinct personalized preference structure or trade-off.
+Then, instead of conditioning the LLM on raw numerical values, we use a symbolic token, <max_utility_index>, representing the index of the utility function that best aligns with a desired personalized reward profile.
+This allows the LLM to learn an implicit mapping between the symbolic token and the characteristics of a personalized response, without requiring explicit numerical reasoning.
+The online fine-tuning process further refines the LLM’s ability to generate responses consistent with these diverse personalized utility profiles.
+Figure 1 illustrates the UC-MOA framework.
+Figure 1: An illustration of the personalization mechanism employed by UC-MOA.
+Each user expresses a distinct preference over multiple objectives, such as humor and helpfulness as shown in the figure.
+These preferences are first mapped to a corresponding utility function from a learned library.
+The index of the selected utility function is then encoded as a symbolic token and incorporated into the input prompt.
+This conditioning guides the LLM in generating a response aligned with the target preference profile.
+3.1 Library of Non-Decreasing Utility Functions
+To enable personalization across a wide spectrum of user needs, the first component of UC-MOA constructs a diverse set of candidate utility functions.
+These functions play a central role in guiding policy optimization toward various distributional preferences, allowing the LLM to adapt its behavior to different personalized application scenarios and user requirements.
+We achieve this through two key design choices: (i) parameterizing each candidate utility function using a monotonic neural network, ensuring they can represent complex, non-linear preference structures; and (ii) optimizing a diversity-promoting loss that encourages discrepancies in both the output values and the local sensitivities (gradients) of these functions.
+Together, these elements ensure that the resulting utility library can capture a broad range of distinct and expressive personalized preference profiles.
 Monotonic Neural Network Parameterization.
-We denote each candidate utility function by $g_{\psi_{i}}:\mathcal{Z}\rightarrow\mathbb{R}$ for $i=1,\dots,M$, where $\mathcal{Z}\coloneqq[0,1]^{K}$ represents the normalized input space.
-The network $g_{\psi_{i}}$ is constructed to be monotonic (i.e., non-decreasing) by enforcing non-negativity on its weight matrices and by employing non-decreasing activation functions.
-This design paradigm follows the approaches proposed in convex neural networks Amos et al.
-(2017) and monotonic architectures such as DPMORL Cai et al.
-(2023), which have been shown to approximate any multivariate non-decreasing function with arbitrarily small error Dugas et al.
-(2009).
-Implementation details are provided in Appendix B.
+We denote each parameterized utility function by $g_{\psi_{i}}:\mathcal{Z}\rightarrow\mathbb{R}$ for $i=1,\dots,M$, where $\mathcal{Z}\coloneqq[0,1]^{K}$ represents the normalized $K$-dimensional reward space and $\psi_{i}$ denotes the corresponding parameters.
+The scalar output of $g_{\psi_{i}}(\mathbf{z})$ quantifies the desirability of a reward vector $\mathbf{z}$ under the specific preference structure $i$.
+Each $g_{\psi_{i}}$ is constructed to be monotonically non-decreasing by enforcing non-negative weights and using non-decreasing activations, following [21, 16].
+This allows approximation of any multivariate non-decreasing function, essential for capturing varied personalized preferences.
+Details are in Appendix D.
 Diversity-Promoting Objective.
-To ensure that the candidate functions span a wide range of utility profiles, we introduce a loss function that encourages dissimilarity both in function values and in their gradients.
-For a given candidate $g_{\psi_{i}}$, we first define the value discrepancy loss as
+To ensure the library ${g_{\psi_{i}}}$ spans a wide range of personalized utility profiles, we optimize a diversity loss $L(\psi_{i})$ that encourages dissimilarity both in function values and in their gradients.
+We first define the value discrepancy loss as
 $L_{\mathrm{val}}(\psi_{i})=\min_{j\neq i}\mathbb{E}_{\mathbf{z}\sim\mathcal{U}(\mathcal{Z})}\Bigl{[}g_{\psi_{i}}(\mathbf{z})-g_{\psi_{j}}(\mathbf{z})\Bigr{]}^{2},$ (1)
-where $\mathcal{U}(\mathcal{Z})$ denotes the uniform distribution over $\mathcal{Z}$. Next, to promote differences in the local behavior, we define the gradient discrepancy loss as
-$\begin{split}L_{\mathrm{grad}}&(\psi_{i})=\;\min_{j\neq i}\mathbb{E}_{\begin{subarray}{c}\mathbf{z},\mathbf{z}^{\prime}\sim\mathcal{U}(\mathcal{Z})\\ \mathbf{z}\neq\mathbf{z}^{\prime}\end{subarray}}\\ &\left[\frac{g_{\psi_{i}}(\mathbf{z}^{\prime})-g_{\psi_{i}}(\mathbf{z})}{\|\mathbf{z}^{\prime}-\mathbf{z}\|}-\frac{g_{\psi_{j}}(\mathbf{z}^{\prime})-g_{\psi_{j}}(\mathbf{z})}{\|\mathbf{z}^{\prime}-\mathbf{z}\|}\right]^{2}.\end{split}$ (2)
+where $\mathcal{U}(\mathcal{Z})$ denotes the uniform distribution over $\mathcal{Z}$ and $\psi_{i}$ denotes the parameters of $g_{\psi_{i}}$.
+Next, to promote differences in the local behavior, we define the gradient discrepancy loss as
+$\begin{split}L_{\mathrm{grad}}(\psi_{i})=\;\min_{j\neq i}\mathbb{E}_{\begin{subarray}{c}\mathbf{z},\mathbf{z}^{\prime}\sim\mathcal{U}(\mathcal{Z})\\ \mathbf{z}\neq\mathbf{z}^{\prime}\end{subarray}}\left[\frac{g_{\psi_{i}}(\mathbf{z}^{\prime})-g_{\psi_{i}}(\mathbf{z})}{\|\mathbf{z}^{\prime}-\mathbf{z}\|}-\frac{g_{\psi_{j}}(\mathbf{z}^{\prime})-g_{\psi_{j}}(\mathbf{z})}{\|\mathbf{z}^{\prime}-\mathbf{z}\|}\right]^{2}.\end{split}$ (2)
 The overall diversity loss for the candidate $g_{\psi_{i}}$ is then defined as a convex combination of these two components:
 $L(\psi_{i})=\mu\,L_{\mathrm{val}}(\psi_{i})+(1-\mu)\,L_{\mathrm{grad}}(\psi_{i}),$ (3)
-where the hyperparameter $\mu\in[0,1]$ balances the relative emphasis on value versus gradient discrepancies.
-Minimizing $L(\psi_{i})$ via gradient descent for each $i=1,\dots,M$ encourages the ensemble $\{g_{\psi_{i}}\}_{i=1}^{M}$ to cover a broad spectrum of utility profiles over $\mathcal{Z}$.
-To satisfy the strict monotonicity condition required for ensuring Distributional Pareto-Optimality, we further modify each candidate by incorporating a small linear term.
-Specifically, we define the final utility function as
+This loss encourages dissimilarity in both function values and gradients, ensuring distinct preference representations. The hyperparameter $\mu\in[0,1]$ balances these terms.
+To ensure robust personalized solutions, we enforce the strict monotonicity required for achieving Distributional Pareto-Optimality, and define the final utility functions as
 $\hat{g}_{\psi_{i}}(\mathbf{z})=g_{\psi_{i}}(\mathbf{z})+\frac{\epsilon}{K}\sum_{k=1}^{K}z_{k},$ (4)
-where $\epsilon$ is a small positive constant (e.g., $\epsilon=0.01$ in our experiments).
-The functions $\{\hat{g}_{\psi_{i}}\}_{i=1}^{M}$ are then employed as the strictly increasing utility functions for guiding policy optimization.
-3.2 Utility Index Selection during Training
-To robustly determine the optimal utility index during training, we employ a percentile ranking strategy in both the offline and online phases.
-For every sample with reward vector $\mathbf{z}$, each candidate utility function $\hat{g}_{\psi_{i}}$ produces a utility score.
-We then compute the percentile rank for the $i$-th utility function as
-$\text{percentile}_{i}=\frac{\text{rank of }\hat{g}_{\psi_{i}}(\mathbf{z})}{N},$ (5)
-where $N$ is the total number of samples, and the rank is determined by the position of $\hat{g}_{\psi_{i}}(\mathbf{z})$ in the sorted array of scores for that function.
-The percentile ranking provides a normalized measure of performance for each utility function, mitigating the effects of scale differences and ensuring that selection is robust across diverse reward distributions.
-During offline and online training, the utility index is then selected as the one with the highest percentile rank.
-We present a concrete example in Appendix B.2.
-3.3 Offline Training
-In the offline phase, our approach establishes a robust grounding of the language model in distributional reward signals by embedding an explicit symbolic indicator of the preferred utility function directly into the training prompt.
-For each training sample, denoted by $(x,y)$ with $x$ representing the input prompt and $y$ the corresponding response, we first compute a multi-objective return vector $\mathbf{z}(x,y)$ and then evaluate a suite of strictly increasing candidate utility functions $\{\hat{g}_{\psi_{i}}\}_{i=1}^{M}$ on $\mathbf{z}(x,y)$.
-The scalar utility values $U_{i}=\hat{g}_{\psi_{i}}(\mathbf{z}(x,y))$ for $i=1,\dots,M$ are then obtained, and for each function, we calculate its percentile rank over the training distribution.
-The maximum utility index is determined as the index with the highest percentile:
-$i^{*}=\arg\max_{i\in\{1,\dots,M\}}\left\{\text{percentile}_{i}\right\}.$ (6)
-We augment the original prompt $x$ by appending a specialized token that encodes this index; that is, the modified prompt is given by
-$x^{\prime}=\text{\#\#\# Prompt: }\{x\}\;\texttt{<max\_utility\_index>}+i^{*}.$
-The model is then fine-tuned via a standard cross-entropy loss:
-$\mathcal{L}_{\text{offline}}(\theta)=-\sum_{t}\log\pi_{\theta}(y_{t}\mid x^{\prime},y_{<t}),$
-ensuring that the model learns to internalize the mapping between input contexts and their corresponding optimal distributional reward profiles.
-This training stage leverages supervised techniques exclusively, thereby avoiding the instability often associated with RL-based methods.
-3.4 Online Fine-Tuning
-To enhance the generalizability of the fine-tuned model, we incorporate additional online stages.
-In the online stage, the large language model generates additional candidate responses and we selectively choose those responses that exhibit superior distributional reward characteristics to fine-tune the model.
-Given an input prompt $x$, the model produces a response $y$ whose multi-objective return $\mathbf{z}(y)$ is computed and then evaluated by each candidate utility function, yielding scalar utility values $U_{i}=\hat{g}_{\psi_{i}}(\mathbf{z}(y))$ for $i=1,\dots,M$.
-For each candidate response, we compute the percentile rank of each utility score as described above, and the optimal utility index is selected as the one with the highest percentile as in Equation (5), resulting in an updated input
-$x^{\prime}=\text{\#\#\# Prompt: }\{x\}\;\texttt{<max\_utility\_index>}+i^{*}.$
-The response $y$ and its corresponding modified prompt $x^{\prime}$ are stored in an online training buffer for further model updates.
-Our online training process also incorporates a rejection sampling mechanism to filter out responses that do not exceed predetermined utility thresholds across multiple reward dimensions.
-The selected samples are then used to update the language model via the same reward-conditional supervised objective as in the offline training phase.
+where $\epsilon$ is a small positive constant, set to $0.01$ in our experiments.
+The functions $\{\hat{g}_{\psi_{i}}\}_{i=1}^{M}$ constitute a library of strictly increasing utility functions, each encoding a distinct way to value multi-objective outcomes, which is critical for enabling diverse personalization.
+3.2 Utility Index Selection via Percentile Ranking
+To determine which pre-defined personalized preference profile best matches a given response’s reward characteristics during training, we employ a percentile ranking strategy based on utility functions.
+For each sample with reward vector $\mathbf{z}$, we compute the utility score $u_{i}=\hat{g}_{\psi_{i}}(\mathbf{z})$ for each function $i$.
+The percentile rank for the $i$-th utility function as
+$\text{percentile}j(\mathbf{z})=\frac{\text{rank of }\hat{g}_{\psi_{j}}(\mathbf{z})\text{ w.r.t. its scores on dataset }\mathcal{D}_{\text{ref}}}{|\mathcal{D}_{\text{ref}}|},$ (5)
+where $\mathcal{D}_{\text{ref}}$ is a reference dataset, such as the offline training set or the current online buffer.
+Similar percentile-based shaping is a standard trick in which it makes the update rule invariant to any monotone transformation of the objective and empirically stabilizes learning under widely varying or drifting reward scales [22, 23].
+The selected utility index $i^{*}$ corresponds to the utility function that yields the highest percentile rank.
+An illustrative example is provided in Appendix D.2.
+3.3 Offline Training: Learning to Map Symbolic Tokens to Personalized Styles
+In the offline phase, we leverage a fixed dataset $\mathcal{D}$ of prompt-response pairs $(x,y)$ to teach the LLM to associate symbolic utility tokens with specific personalized response styles.
+For each training sample $(x,y)$, we compute its normalized reward vector $\mathbf{z}(x,y)\in[0,1]^{K}$.
+Using percentile ranking (Equation (5) relative to $\mathcal{D}$), we identify the index $i^{*}$ of the utility function $g_{\psi_{i^{*}}}$ for which $y$ achieves the highest relative performance, indicating that $y$ best embodies the preference structure $i^{*}$.
+The prompt $x$ is augmented with a symbolic token representing this optimal personalized preference index:
+$x^{\prime}=\text{\#\#\# Prompt: }\{x\}\;\texttt{<max\_utility\_index>}+i^{*}.$ (6)
+The model $\pi_{\theta}$ is trained via standard supervised fine-tuning (SFT) to predict the original response $y$ given the augmented prompt $x^{\prime}$, using a cross-entropy loss:
+$\mathcal{L}_{\text{offline}}(\theta)=-\mathbb{E}_{(x,y)\sim\mathcal{D}}\left[\sum_{t=1}^{|y|}\log\pi_{\theta}(y_{t}|x^{\prime},y_{<t})\right]$ (7)
+Crucially, the LLM learns an implicit mapping: it associates the discrete token (e.g., <max_utility_index> 5) with the linguistic characteristics of responses optimal for the underlying personalized utility profile (e.g., $\hat{g}_{\psi_{5}}$).
+This avoids reliance on numerical interpretation, enabling more reliable personalization by learning a policy $\pi_{\theta}(y|x,i^{*})$ conditioned on the latent user personalization structure $i^{*}$.
+3.4 Online Fine-Tuning: Refining Personalized Response Generation
+Online fine-tuning allows the LLM to actively generate responses and iteratively refine its ability to align with diverse personalized preference profiles.
+Given an input prompt $x$, the model $\pi_{\theta}$ generates a response $y$.
+We compute its reward vector $\mathbf{z}(y)$ and evaluate it with all utility functions ${\hat{g}_{\psi_{i}}}$.
+The optimal utility index $i^{*}$ is determined via percentile ranking (Equation (5) relative to the current online buffer), identifying the preference profile the response best satisfies.
+We further form an augmented prompt $x^{\prime}$ using $i^{*}$ as in Equation (6).
+High-quality pairs through rejection sampling are added to an online buffer and used to update $\pi_{\theta}$ via Equation (7).
+This online loop enables the model to explore beyond the initial dataset and improve its capability to generate outputs tailored to a wide range of personalized utilities.
 3.5 Inference Stage
-At inference time, the user specifies a preference vector $\mathbf{w}=[w_{1},\ldots,w_{N}]$ that encodes the relative importance of each reward dimension $r_{1},\ldots,r_{N}$.
-To translate these preferences into a concrete reward profile, we apply a linear preference-to-reward mapping for each dimension:
-$f_{i}(w_{i})=w_{i}\times(r_{i}^{\mathrm{max}}-r_{i}^{\mathrm{min}})+r_{i}^{\mathrm{min}},$
-where $r_{i}^{\mathrm{max}}$ and $r_{i}^{\mathrm{min}}$ denote the maximum and minimum values of the $i$-th reward observed in the training dataset.
-This mapping yields a target reward vector $\mathbf{z}=[f_{1}(w_{1}),\dots,f_{N}(w_{N})]$ that accurately reflects the user’s preferences while normalizing the reward scales across dimensions.
-Subsequently, the target reward vector $\mathbf{z}$ is input into each of the pre-trained, strictly increasing utility functions $\{\hat{g}_{\psi_{i}}\}_{i=1}^{M}$ to obtain scalar utility values $U_{i}=\hat{g}_{\psi_{i}}(\mathbf{z})$ for $i=1,\dots,M$.
-at inference time the reward distribution is more stable and the set of candidate responses is smaller; therefore, we directly use the argmax operation to select the index corresponding to the maximum expected utility:
-$i^{*}=\arg\max_{i\in\{1,\dots,M\}}U_{i}.$ (7)
-This index $i^{*}$ is encoded into the input prompt via a symbolic token—specifically, the prompt is augmented as
-$x^{\prime}=\text{\#\#\# Prompt: }\{x\}\;\texttt{<max\_utility\_index>}+i^{*},$
-thereby conditioning the language model to generate responses that are robustly aligned with the user’s specified multi-objective reward preferences while obviating the need for direct numerical processing.
+At inference time, UC-MOA offers flexible mechanisms for users to articulate their personalized multi-objective preferences, ensuring the LLM generates responses tailored to their specific needs.
+The user can specify their preferences in one of two ways:
+Method 1: Direct Preference Vector Specification.
+Similar to the previous description, the user can provide a direct preference vector $\mathbf{w}=[w_{1},\ldots,w_{K}]$, where each $w_{j}$ indicates the relative importance or desired level for the $j$-th reward dimension.
+This vector $\mathbf{w}$ is transformed into a target reward vector $\mathbf{z}_{\text{target}}$ (e.g., via linear mapping normalized by observed reward ranges from the training data).
+This $\mathbf{z}_{\text{target}}$ represents the user’s ideal outcome in the multi-objective reward space.
+We then evaluate $\mathbf{z}_{\text{target}}$ using all pre-trained, strictly increasing utility functions $\{g_{\psi_{j}}\}_{j=1}^{M}$ from our library.
+The index $i^{*}$ of the utility function that best aligns with this explicit preference is selected by finding the function that maximizes the utility for $\mathbf{z}_{\text{target}}$:
+$i^{*}=\arg\max_{j\in\{1,\dots,M\}}g_{\psi_{j}}(\mathbf{z}_{\text{target}}).$ (8)
+This $i^{*}$ identifies the pre-trained personalized utility profile that most closely matches the user’s explicitly stated preference trade-offs.
+Method 2: Preference Elicitation from Labeled Examples.
+Alternatively, if a user finds it more intuitive to demonstrate their preferences through examples rather than specifying a numerical vector, they can provide a small set of $L$ labeled examples, $\mathcal{D}_{\text{user}}=\{(x_{l},y_{l},s_{l})\}_{l=1}^{L}$.
+Each example consists of a prompt $x_{l}$, a corresponding response $y_{l}$ (which could be generated by a baseline model or written by the user), and a scalar preference score $s_{l}$ assigned by the user to that response, reflecting how well $y_{l}$ meets their personalized criteria for that specific prompt.
+For each example $(x_{l},y_{l},s_{l})$ in $\mathcal{D}_{\text{user}}$:
+  1. 1.
+    We first compute the multi-dimensional normalized reward vector $\mathbf{z}(x_{l},y_{l})\in[0,1]^{K}$ for the response $y_{l}$.
+  2. 2.
+    Then, for each pre-trained utility function $\hat{g}_{\psi_{j}}$ in our library, we calculate its output score $\hat{g}_{\psi_{j}}(\mathbf{z}(x_{l},y_{l}))$.
+To identify the utility function from our library that best represents the user’s demonstrated personalized preferences, we select the index $i^{*}$ that minimizes the Mean Squared Error (MSE) between its outputs and the user’s provided scores across all labeled examples:
+$i^{*}=\arg\min_{j\in\{1,\dots,M\}}\frac{1}{L}\sum_{l=1}^{L}\left(g_{\psi_{j}}(\mathbf{z}(x_{l},y_{l}))-s_{l}\right)^{2}.$ (9)
+This $i^{*}$ corresponds to the pre-trained personalized utility profile whose valuation of responses most closely mimics the user’s scoring behavior on their provided examples.
+This method allows for personalization even when users cannot easily articulate their preferences as a numerical vector.
+Generating the Personalized Response.
+Once the optimal utility index $i^{*}$ has been determined (either via Equation (8) or Equation (9)), it signifies the personalized preference structure the LLM should adopt.
+The original input prompt $x_{\text{new}}$ (for which the user desires a personalized response) is augmented using the symbolic token format:
+$x^{\prime}_{\text{new}}=\text{\#\#\# Prompt: }\{x_{\text{new}}\}\;\texttt{<max\_utility\_index>}+i^{*}.$ (10)
+The fine-tuned LLM $\pi_{\theta}$ then generates a response conditioned on $x^{\prime}_{\text{new}}$.
+This response is tailored to the user’s specified or demonstrated personalized multi-objective preferences, as captured by the selected utility function $g_{\psi_{i^{*}}}$, without requiring the LLM to directly process raw numerical reward targets or complex preference descriptions at inference time.
+3.6 Theoretical Guarantees
+UC-MOA’s symbolic conditioning not only avoids numerical reasoning pitfalls but also provides theoretical underpinnings for achieving personalized alignment that is distributionally Pareto-optimal.
+Let $\mathcal{Z}\!=\![0,1]^{K}$ be the normalized $K$-dimensional reward space and $\mathcal{U}^{\uparrow}$ represent the set of strictly increasing utility functions $g_{\psi_{i}}:\mathcal{Z}\!\rightarrow\!\mathbb{R}$.
+Given a prompt $x$ and a utility index $i\!\in\!\{1,\dots,M\}$, our fine-tuned model $\pi_{\theta}$ exposes a utility-conditioned sub-policy
+$\pi_{\theta,i}(\,\cdot)\;=\;\pi_{\theta}\!\bigl{(}\,\cdot\mid x,\langle\texttt{max\_utility\_index}\rangle+i\bigr{)}.$
+First, we have the definition of the expected utility:
+Definition 1 (Expected utility).
+For a policy $\pi$, let $\mu_{\pi}$ denote the distribution of return vectors. Then we can define the expected utility of $\pi$ under the utility function $g_{\psi_{i}}$ as
+$J_{g_{\psi_{i}}}(\pi)\;=\;\mathbb{E}_{\mathbf{z}\sim\mu_{\pi}}\bigl{[}g_{\psi_{i}}(\mathbf{z})\bigr{]}.$
+Then, we hold the following assumption:
+Assumption 1 (Utility-optimal sub-policies for Personalization).
+For every $i\!\in\!\{1,\dots,M\}$, $\pi_{\theta,i}\in\arg\max_{\pi}J_{g_{\psi_{i}}}(\pi),$ i.e.
+each sub-policy maximizes the expected value of its associated strictly increasing utility $g_{\psi_{i}}$.
+Finally, we present the following theorem:
+Theorem 1 (Utility-Conditional Distributional Pareto-Optimality).
+Under Assumption 1, for any generated response $y\!\sim\!\pi_{\theta,i}(x)$ with reward vector $\mathbf{z}(y)$, let
+$i^{\star}\;=\;\arg\max_{j\in\{1,\dots,M\}}g_{\psi_{j}}\!\bigl{(}\mathbf{z}(y)\bigr{)}.$
+Then the sub-policy $\pi_{\theta,i^{\star}}$ is distributionally Pareto-optimal.
+A policy being distributionally Pareto-optimal means no other policy stochastically dominates it on all reward dimensions while being strictly better on at least one.
+Theorem 1 (proof in Appendix F) ensures that by selecting among sub-policies optimized for diverse, strictly increasing utility functions (each representing a distinct personalized preference profile), UC-MOA achieves personalized responses that are also distributionally Pareto-optimal.
+This result grounds UC-MOA’s symbolic token mechanism in a rigorous optimality guarantee: every utility index corresponds to a distinct, nondominated, and distributional trade-off in the multi-objective reward space.
 4 Evaluation
-In this section, we start with the experimental setup, including the tasks, baselines, and implementation.
-We then present our experiment design, followed by experiment results and analysis.
+This section details the empirical evaluation of our proposed method.
+We first describe the experimental setup, including the tasks, baselines, and implementation details.
+Subsequently, we present a comprehensive assessment of our method’s effectiveness.
+This is primarily demonstrated through a user study designed to capture qualitative preference alignment, and further substantiated by quantitative metrics focusing on inference consistency and Pareto front optimality .
+Finally, we analyze the computational efficiency of our approach.
+More experimental results can be found in Appendix E.
 We provide our code in the anonymous link https://anonymous.4open.science/r/UC-MOA-8D36.
@@
 Tasks.
-Following RiC (Yang et al., 2024c), we select two text generation tasks for our experiments, i.e., Helpful Assistant task (Bai et al., 2022) and the Reddit Summary task (Stiennon et al., 2020).
+Following RiC [13], we select two text generation tasks for our experiments, i.e., Helpful Assistant task [2] and the Reddit Summary task [3].
 For the Helpful Assistant task, we utilize three open-sourced reward models on Huggingface to measure three different objectives “harmless”, “helpful”, and “humor”.
@@
 We focus on three multi-objective subtasks: (1) “summary” vs “faithful”; (2) “hallucinated” vs “faithful”; (3) “summary” vs “hallucinated”.
-We provide more details about the datasets and the reward models in Appendix A.
-Baseline methods.We compare our method with two state-of-the-art multi-objective RLHF methods, Rewarded Soups, and Reward-in-Context (RiC).
-We also report the performance of the base model and the supervised fine-tuned (SFT) model.
+More details on datasets and reward models are in Appendix C.
+Baseline methods.
+We compare our method with three state-of-the-art multi-objective alignment methods for personalization, i.e., Multi-Objective Reinforcement Learning from Human Feedback (MORLHF), Rewarded Soups, and Reward-in-Context (RiC).
+We also report the performance of the base LLaMA 2 7B model and the supervised fine-tuned (SFT) model.
+We document the implementation of the baseline methods in Appendix D.3.
 Implementation.
@@
 We train the machine learning models on a server with 8 NVIDIA A100 80GB GPUs and 4TB memory for all methods.
-We select the LLaMA 2 7B-hf model as the base model for alignment following the setting of Yang et al.
-(2024c).
-We list the implementation details (e.g., network structure for the parameterized utility function, hyper-parameter choices) in Appendix B.
-4.2 Experiment Design
-We use the following experiments to evaluate the effectiveness and efficiency of our method, and other factors that influence the performance (e.g., alternative design choices, hyper-parameters).
-Experiment I.
-To evaluate the effectiveness of our method, we compare the curves of the multi-dimensional average test rewards that correspond to the empirical Pareto fronts generated by different methods.
-A method demonstrating an outer curve indicates better performance, as it captures a wider range of trade-offs across objectives.
-We plot the Pareto fronts of all baseline methods across the Helpful Assistant task and the Reddit Summary task.
-Experiment II.
-We analyze the computational efficiency of each method by recording the training time of each baseline method.
-We run experiments with NVIDIA A100 80GB GPUs and report the GPU hours utilized by each method.
-Experiment III.
-To validate the necessity of the online stage, we perform an ablation study to compare the performance of pure offline training (“w/o online”), a single online iteration (“w/ online iter 1”), and two online iterations (“w/ online iter 2”).
-Experiment IV.
-We assess the consistency of our method during the inference phase by measuring the alignment between the user-specified target objective and the reward outcomes obtained from the LLM’s generated responses.
-Each prompt provided to the LLM embeds a target objective requirement (e.g., in the form <max_utility_index> a, where the letter encodes the desired objective).
-During evaluation, the LLM is tasked with generating responses that adhere to these preset requirements.
+We select the LLaMA 2 7B-hf model as the base model for alignment following the setting of [13].
+We list the implementation details of our method (e.g., network structure for the parameterized utility function, hyper-parameter choices) in Appendix D.
+4.2 Evaluating Effectiveness
+We evaluate the effectiveness of our method through both qualitative user assessments and a suite of quantitative metrics.
+4.2.1 User Study: Qualitative Preference Alignment
+Design.
+To directly assess how well the models satisfy genuine user preferences, we perform a single-blind A/B test on the humorhelpfulness trade-off.
+We first curate five diverse dialogue prompts that explicitly ask the assistant to be either funny or helpful.
+For each prompt, we generate two candidate answers: one with the state-of-the-art method RiC and one with our method UC-MOA.
+The order of the two answers is randomly shuffled and shown side-by-side to twelve independent volunteer raters, none of whom are authors and all of whom were blind to system identity.
+Each rater casts one vote per prompt for the response that better satisfies both objectives, yielding a total of $5\times 12=60$ binary votes.
+The prompts, outputs, and anonymized comments are included in Appendix G.
+Our user study has an IRB approval.
+Results.
+UC-MOA was preferred in $46$ out of the $60$ comparisons (76.7%), whereas RiC was preferred in $14$ (23.3%).
+A two-sided binomial test rejects the null hypothesis that the two systems are equally likely to be chosen ($p<0.001$).
+Qualitatively, raters commented that UC-MOA answers felt more on-point and balanced comedy with solid advice, whereas RiC outputs were occasionally funny but off–topic or helpful but bland.
+These findings corroborate the quantitative superiority of UC-MOA and highlight its ability to internalize nuanced, non-linear user utilities.
+4.2.2 Quantitative Evaluation of Effectiveness
+We further assess our method’s effectiveness using several quantitative measures.
+Consistency in Inference Alignment.
+To evaluate UC-MOA’s ability to adhere to specified preferences during inference, we embed target utility indices (<max_utility_index> i) into prompts and measure whether the generated responses’ reward profiles align with the targeted utility function.
 For each generated response, we compute a set of utility values using a suite of ten reward models.
@@
 We then quantify consistency by comparing the expected utility and the value distribution of each utility function with the target objective extracted from the prompt.
-Experiment V.
-To investigate how the number of utility functions impacts the performance of our method, we vary the number of utility functions from {5,10,15} and test the performance of our method with different numbers of utility functions.
-Experiment VI.
-We investigate an alternative design choice of the utility functions, i.e., linear functions.
-The linear function is designed in the form of $w^{T}x$ where $||w||_{2}=1$.
-To construct $N$ linear utility functions, we uniformly sample $N$ points from the unit circle and set $w$ accordingly.
-We plot the Pareto fronts using linear utility functions and our method across all tasks.
-Experiment VII.
-We also investigate two metrics developed specifically for assessing distributional preference, i.e., Constraint Satisfaction and Variance Objective as in previous distributional studies (Cai et al., 2023).
-They are designed to underscore the optimality of multivariate distributions associated with LLMs.
-In terms of Constraint Satisfaction, we randomly generate $M=100$ constraints.
-The Constraint Satisfaction metric is then computed by considering the highest probability within the policy set that satisfies each individual constraint.
-The Variance Objective metric, on the other hand, involves generating $M=100$ random linear weights.
-These weights are applied to both the expected returns and the standard deviations of returns in each dimension.
-This objective encourages attaining greater expected returns while simultaneously increasing variance, thereby catering to dynamic preferences.
-Further details about the implementation of these evaluation metrics are provided in Appendix B.4.
-4.3 Experiment Results.
-Performance comparison.
-We compare the multi-dimensional average test rewards obtained from the empirical Pareto fronts generated by different methods on two distinct tasks: the Helpful Assistant task and the Reddit Summary task.
-We show the results of “harmless” vs.
-“humor” task and “deberta” vs.
-“faithful” in Figure 2 while leaving the rest of the results in Appendix C.1.
-We observe that our method consistently produces an outer Pareto curve that dominates those of Rewarded Soups and RiC, indicating its superior ability to capture a broader spectrum of trade-offs.
+In a case study on the “harmless” vs.
+“humor” task, we appended the requirement “<max_utility_idx> i” to each prompt to encourage the model to follow the first preference (i.e., maximizing the 9th utility function).
+We then used reward models to evaluate the generated responses.
+A detailed analysis in Appendix E.1 presents the distribution of normalized utility values for each utility function over our evaluation data.
+Notably, the 9th utility function achieves the highest mean normalized value, indicating that the fine-tuned model aligns its responses with the specified preference requirement.
+This demonstrates the model’s ability to selectively prioritize and adhere to user-designated objectives.
+Pareto Front Analysis.
+We compare the empirical Pareto fronts generated by each method based on multi-dimensional average test rewards.
+A method achieving an outer curve demonstrates superior effectiveness in capturing diverse, high-quality trade-offs between objectives.
+We plot these fronts for all methods across all subtasks.
+Figure 2 shows the results for the “harmless” vs.
+“humor” and “deberta” vs.
+“faithful” tasks (additional results in Appendix E.2).
+We observe that our method consistently produces an outer Pareto curve that dominates those of MORLHF, Rewarded Soups, and RiC, potentially reflecting its ability to handle complex, non-linear preferences via learned utility functions.
 Figure 2: Pareto fronts of “harmless” vs “humor” and “deberta” vs.
@@
 Each marker on the plotted curves represents the average test rewards corresponding to a particular user preference.
+Note that a curve that covers a larger area indicates a better performance of the corresponding method.
+4.3 Efficiency Analysis
 Computational Cost.
-We measure the time complexity of all methods in the two-objective Helpful Assistant task and the two-objective Reddit Summary task using NVIDIA A100 80GB GPUs.
-We report the time cost (in the form of GPU hours) in Table 1 and Table 2.
-We observe that Rewarded Soups consistently consumes the most GPU hours across both tasks, while our method (w/ online iter2) achieves notably lower costs.
-Meanwhile, RiC requires more computing time than our method, though it remains more cost-friendly than Rewarded Soups.
-These findings demonstrate that our approach not only delivers strong performance (as shown in previous experiments) but also offers a clear advantage in terms of computational efficiency.
-  Method GPU hours
-  Rewarded Soups 69.5
-RiC 20.3
-Ours w/o online 8.9
-Ours w/ online iter1 10.1
-Ours w/ online iter2 11.3
-Table 1: Comparison of GPU hours in the Helpful Assistant task, where the number of preference $M=10$, number of rewards $N=2$.
-  Method GPU hours
-  Rewarded Soups 15.9
-RiC 5.7
-Ours w/o online 2.0
-Ours w/ online iter1 2.4
-Ours w/ online iter2 2.8
-Table 2: Comparison of GPU hours in the Reddit Summary task, where the number of preference $M=10$, number of rewards $N=2$.
-Necessity of online stage.
-We evaluate the impact of the online stage by comparing models trained purely offline (“w/o online”) with those undergoing one (“w/ online iter 1”) and two online iterations (“w/ online iter 2”) across all tasks.
-We present the performance results for “harmless” vs.
-“humor” and “summary” vs.
-“faithful” tasks in Figure 3 while leaving the rest of the results in Appendix C.2.
-We observe that models incorporating online iterations consistently outperform the offline-only baseline.
-Notably, even a single online iteration yields significant improvements, with further performance gains observed when two iterations are applied, thereby underscoring the necessity of the online stage in our training framework.
-The online stage is important because it enables the model to generalize over unseen samples in the offline dataset, and ultimately enhancing its ability to generate appropriate responses corresponding to the required utility function index.
-Figure 3: Ablation study of the online stage on “harmless” vs. “humor” task and “summary” vs. “faithful” task.
-Consistency.
-We conduct a case study on the “harmless” vs.
-“humor” task, appending the requirement “<max_utility_idx> i” to each prompt to encourage the model to follow the first preference (i.e., maximizing 9th utility function).
-We then use reward models to evaluate the generated responses.
-We provide a detailed analysis in Appendix C.3 to present the distribution of normalized utility values for each utility function over our evaluation data.
-Notably, the 9th utility function achieves the highest mean normalized value, indicating that the fine-tuned model aligns its responses with the specified preference requirement.
-This demonstrates the model’s ability to selectively prioritize and adhere to user-designated objectives.
-Number of utility functions.
-We vary the number of utility functions over the set 5, 10, 15 and compare the resulting empirical Pareto fronts in Figure 4.
-We leave the results of the remaining four tasks in Appendix C.4.
-With 5 utility functions, the generated Pareto front is slightly inferior to those produced with 10 and 15 utility functions.
-Specifically, the front with 5 utility functions captures a narrower range of optimal trade-offs across the objectives.
-However, when the number of utility functions is increased to 10 or 15, the resulting Pareto fronts exhibit comparable performance.
-Increasing the number of utility functions beyond 10 does not yield substantial gains.
-These results demonstrate that our method is insensitive to the precise number of utility functions, provided that a sufficient number (around 10) is used to effectively cover the preference space.
-In practical terms, employing 10 utility functions strikes a good balance between performance and computational efficiency, as further increases to 15 do not produce notable improvements.
-Figure 4: Ablation study of different numbers of utility functions on “humor” vs. “helpful” and “deberta” vs. “faithful” tasks.
-Variants of utility functions.
-We explored the impact of using linear utility functions, designed as $w^{T}x$ where $||w||_{2}=1$, and compared the performance with our original approach across all tasks.
-The generated Pareto fronts, as shown in Figure 5 and Appendix C.5, illustrate that our method continues to deliver strong results even with linear utility functions.
-However, there is still a performance gap between the results achieved with distributional Pareto-optimal utility functions and those obtained using linear functions.
-While linear functions still produce reasonably good trade-offs, they fail to capture the full range of optimal solutions attainable by our method, which leverages distributional utility representations.
-This highlights the advantage of our approach in handling more complex utility functions, demonstrating its superior flexibility in navigating multi-objective optimization.
-Figure 5: Ablation study of using linear utility functions.
-Constraint satisfaction and variance objective.
-In our case study on the “harmless” vs.
-“humor” task, we use 2,000 inference samples and compute both metrics for our method, Rewarded Soup, and RiC.
-As summarized in Table 3, our method achieves the highest overall constraint satisfaction ratio and exhibits the largest variance objective value, outperforming both Rewarded Soup and RiC.
-This indicates that our approach offers a more comprehensive distributional coverage, effectively handling diverse user preferences.
-  Method Constraint Satisfaction Variance
-RiC 0.5637 0.0662
-Rewarded Soups 0.6744 0.1631
-Ours 0.7184 0.5160
-Table 3: Comparison of Methods for Constraint satisfaction and variance objective.
+We measure and compare the training time (GPU hours on 80GB A100) required by each method to achieve its reported performance, providing insight into practical feasibility.
+Table 1 reports the computational cost.
+As expected, methods requiring multiple model training or extensive tuning, like Rewarded Soups and MORLHF, exhibit significantly higher costs.
+RiC is more efficient than these but still requires considerably more time than our approach.
+UC-MOA, even with two online iterations, demonstrates superior computational efficiency compared to all other multi-objective baselines (MORLHF, Rewarded Soups, and RiC), offering a compelling balance between strong performance and resource usage.
+Table 1: Comparison of GPU hours for different methods on Helpful Assistant and Reddit Summary tasks, where the number of preference $M=10$, number of rewards $N=2$.
+  Method Helpful Assistant Reddit Summary
+  MORLHF 726.5 216.9
+Rewarded Soups 69.5 15.9
+RiC 20.3 5.7
+Ours w/o online 8.9 2.0
+Ours w/ online iter1 10.1 2.4
+Ours w/ online iter2 11.3 2.8
+We also conduct an ablation study to investigate how other factors (i.e., the necessity of the online stage, number of utility functions, and variants of utility functions) influence the performance of our method.
+The results are presented in Appendix E.3.
+5 Limitations
+We acknowledge several limitations in the proposed UC-MOA framework.
+First, the overall accuracy of this framework is sensitive to the quality of the multi-objective reward models used to generate the reward vectors $\mathbf{z}$.
+Noise or systematic biases in these underlying reward signals can inevitably propagate into the utility calculations and affect the selection of the conditioning index $i^{*}$, potentially leading to suboptimal or misaligned personalization.
+Second, the initial training of the diverse utility library can be computationally intensive, with complexity scaling in proportion to the number of utility functions and reward dimensions.
+Finally, although UC-MOA avoids embedding sensitive preference details directly into model parameters by using symbolic indices, practical deployment should still address privacy risks associated with collecting and processing user preferences.
+In addition, fairness concerns must be carefully addressed, including the potential reinforcement of existing biases and the creation of echo chambers resulting from fine-grained personalization.
 6 Conclusion
-We propose Utility-Conditioned Multi-Objective Alignment (UC-MOA), a novel multi-objective LLM alignment framework.
-By transforming user-specified preference into symbolic tokens through a diverse ensemble of strictly increasing, non-linear utility functions, UC-MOA circumvents the numerical sensitivity issues seen in methods like RiC while avoiding the high computational costs of approaches such as MORLHF, Rewarded Soups, and MODPO.
-Our experiments demonstrate that UC-MOA achieves superior Pareto fronts across complex reward dimensions, evidencing robust and scalable alignment with human values.
-Limitations
-Representation of Human Preferences.
-Although UC-MOA effectively captures the distributional properties of human feedback, its performance heavily depends on the quality and diversity of the underlying preference data.
-Real-world human feedback is often noisy or biased, and addressing such discrepancies requires robust preprocessing and annotation methods.
-Additionally, certain complex preferences might not be adequately represented by the utility functions, potentially limiting the model’s applicability in diverse contexts, such as safety-critical domains like healthcare and autonomous systems.
-Computational and Scaling Challenges.
-The proposed framework improves efficiency compared to existing methods, but it still requires substantial computational resources when scaling to tasks with numerous objectives or high-dimensional reward spaces.
-The reliance on large-scale GPUs and the requirement for iterative online updates may limit its adoption in settings with constrained computational budgets.
-Further optimization of the model architecture and training process could help mitigate these challenges, making the framework more accessible and enabling faster experimentation.
-Generalizability Across Tasks.
-While UC-MOA has demonstrated superior alignment performance in text generation and summary tasks, its applicability to other domains, such as multimodal tasks, remains unexplored.
-Extending the framework to these areas might introduce additional challenges, including the need to handle dynamic environments and integrate feedback in near real-time.
-Adapting and validating this approach across a broader range of use cases to evaluate its generalizability and robustness is a potential direction for future research.
-Ethical Considerations
-While multi-objective optimization and reward modeling can enhance the alignment of large language models, our method must be applied responsibly to avoid harmful outcomes.
-The potential biases or harmfulness in training data or reward models may mislead the final LLM’s behavior.
-Careful curation of training and evaluation datasets, combined with auditing of model outputs, is essential to mitigate these risks.
-Appendix A Details of Datasets
+We present Utility-Conditioned Multi-Objective Alignment (UC-MOA), a novel framework for effective multi-objective personalization of Large Language Models (LLMs).
+UC-MOA leverages a diverse library of non-linear utility functions to represent distinct personalized preference structures, conditioning the LLM via symbolic tokens rather than direct numerical inputs.
+This “utility-conditioned” approach circumvents the numerical sensitivity issues and high computational costs of prior methods, enabling the LLM to learn implicit mappings to varied personalized response styles.
+Grounded in the principle of distributional Pareto-optimality, our experiments, including a user study, demonstrate UC-MOA achieves superior Pareto fronts and better satisfy diverse user preferences, offering a scalable solution for personalized and aligned LLM interactions.
+Appendix A Boarder Impact
+The advancement of multi-objective personalization frameworks like UC-MOA offers substantial positive societal potential by enabling Large Language Models (LLMs) to align more closely with diverse individual user needs.
+This can enhance the utility and user-centricity of AI across various applications, including adaptive educational systems, personalized accessibility tools, and more nuanced creative and informational assistants.
+By facilitating alignment with complex, non-linear preference structures, such technologies can transform AI interactions from generic exchanges to more individually beneficial and collaborative engagements.
+Conversely, sophisticated LLM personalization introduces significant societal risks.
+The capacity for fine-grained adaptation could be exploited for malicious purposes, such as generating targeted disinformation, creating highly convincing fraudulent profiles, or deploying manipulative social bots.
+Fairness considerations are paramount, as personalization mechanisms might inadvertently perpetuate or amplify algorithmic bias if certain preference profiles correlate with demographic groups, potentially leading to disparate impacts or the creation of strong filter bubbles.
+Furthermore, the collection and processing of user preference data, essential for personalization, raise substantial privacy concerns.
+Even with correct functioning, errors in preference inference or alignment could lead to unintended harms.
+Mitigation strategies, including robust user controls, privacy-preserving techniques, bias auditing, and careful consideration of deployment contexts, are imperative to responsibly harness the benefits of personalized LLMs while safeguarding against their potential negative consequences.
+Appendix B Safeguards
+The responsible deployment of UC-MOA necessitates several safeguards.
+Models developed using this framework should be subject to controlled release, incorporating usage guidelines and robust safety filters.
+The integrity of datasets for training utility functions and LLMs is paramount, requiring meticulous curation to mitigate bias and remove unsafe content.
+User preference data must be managed under stringent privacy protocols, ensuring transparency and consent.
+Regular auditing of learned utility functions and underlying reward models for biases is essential for fairness.
+Furthermore, continuous monitoring of system behavior, particularly across different personalized utility profiles, is crucial for detecting and addressing potential misuse or emergent harms.
+These measures are integral to fostering the responsible application of personalized LLM technologies.
+Appendix C Details of Datasets
 In this section, we introduce the HH-RLHF dataset and Reddit Summary dataset in more detail.
 We provide the links to the HH-RLHF dataset and the three reward models we used for the experiments in Table 2.
+Table 2: Datasets and Reward Models.
 HH-RLHF Dataset
@@
 Faithful Reward bart-faithful-summary-detector
-Table 4: Datasets and Reward Models.
 HH-RLHF Dataset
-The HH-RLHF dataset, as described in Bai et al.
-(2022), was collected to support research on aligning language models using human preferences.
+The HH-RLHF dataset, as described in [2], was collected to support research on aligning language models using human preferences.
 The dataset comprises approximately 160,000 prompts each paired with two responses generated by a language model.
@@
 This dual-annotation design allows for training reward models that capture subtle aspects of human judgment.
-Note that we follow Yang et al.
-(2024d) and use “chosen” data to construct our offline dataset.
+Note that we follow [20] and use “chosen” data to construct our offline dataset.
 Reddit Summary Dataset
-The Reddit Summary dataset, introduced in Stiennon et al.
-(2020), is intended for research in abstractive summarization with human feedback.
+The Reddit Summary dataset, introduced in [3], is intended for research in abstractive summarization with human feedback.
 The dataset consists of approximately 14,900 Reddit posts paired with corresponding summaries.
 Each entry comprises a post and its summary, where the summary is either the original TL;DR provided by the poster or a human-annotated summary aimed at capturing the post’s main ideas.
-Appendix B Implementation Details
-B.1 Utility Functions
+Appendix D Implementation Details
+D.1 Utility Functions
 Non-Decreasing Neural Network.
@@
 In the implementation, the return $\mathbf{z}_{t}$ is normalized before feeding into the utility function to compute the scalar reward.
-B.2 Prompt Labeling
+D.2 Prompt Labeling
 We provide an example of our prompt labeling technique in the “harmless” vs.
@@
 Figure 3: An example of our prompt labeling technique in “harmless” vs. “helpful” dataset.
-B.3 Hyperparameter Choices
-Table 5 summarizes the hyper-parameter setting for our experiments.
+D.3 Baseline Implementations
+Regarding baseline approaches (MORLHF, Rewarded Soups, and RiC), we use the code released by the authors of RiC from https://github.com/YangRui2015/RiC.
+D.4 Hyperparameter Choices
+Table 3 summarizes the hyper-parameter setting for our experiments. In our implementation, we choose the threshold as 0.7 for percentile ranking.
+Table 3: Hyper-parameter settings in our experiments.
 Training Phase
@@
 do_sample: True
-Table 5: Hyper-parameter settings in our experiments.
-B.4 Distributional Metrics
-We use two novel metrics as in Cai et al. (2023) for assessing distributional preference: Constraint Satisfaction and Variance Objective.
-For Constraint Satisfaction, we define $u(p,\pi)$ to be the probability of return random variance $z\sim\mathbf{Z}(\pi)$ that satisfies the constraint $p$, representing how a policy satisfies the given constraint.
-We randomly generate each constraint defined by $n$ linear constraints represented by $\{\mathbf{z}|\forall i,\mathbf{w}_{i}^{T}\mathbf{z}\geq c_{i}\}$, where $w_{i}$ is uniformly sampled from the surface $\mathbf{w}_{i}\succeq 0,\mathbf{1}^{T}\mathbf{w}_{i}=1$, and $c_{i}$ is sampled from the distribution of $\mathbf{U}(\min_{\mathbf{z}\in\mathbf{Z}}\mathbf{w}_{i}^{T}\mathbf{z},\max_{\mathbf{z}\in\mathbf{Z}}\mathbf{w}_{i}^{T}\mathbf{z})$, where $\mathbf{Z}$ is the set of all return samples of the final policies learned by DPMORL.
-For Variance Objective, we define $u(p,\pi)$ to be a weighted sum of expected return $\mathbb{E}_{\mathbf{z}\sim\mathbf{\mu}(\pi)}\mathbf{z}$ and standard deviation of return $\sqrt{\text{Var}_{\mathbf{z}\sim\mathbf{\mu}(\pi)}[\mathbf{z}]}$.
-We randomly generate $\mathbf{w}_{i}\in\mathbb{R}^{2K}$, where $\mathbf{w}_{i}\succeq 0$ and $\mathbf{1}^{T}\mathbf{w}_{i}=1$.
-We let $\mathbf{w}_{i}=[\mathbf{w}_{i}^{1},\mathbf{w}_{i}^{2}]$, where $\mathbf{w}_{i}^{1},\mathbf{w}_{i}^{2}\in\mathbb{R}^{K}$, and use $\mathbf{w}_{i}^{1}$ and -$\mathbf{w}_{i}^{2}$ as the weight for expected return and standard derivation of return respectively.
-For both of the metrics, we define $u(p,\pi)$ to be the satisfaction score of policy $\pi$ under preference $p$.
-We generate $M=100$ number of preference $p_{1},p_{2},\cdots,p_{M}$.
-For both of the Constraint Satisfaction and Variance Objective metric, the mean score is defined as
-$\frac{1}{M}\sum_{i=1}^{M}\max_{j\in\{1,2,\cdots,N\}}u(p_{i},\pi_{j})$
-The mean score in the above equation measures the best performance of satisfying the preference of the user, averaged in $M=100$ preferences.
-A higher mean score represents a better ability to meet different preferences.
-Appendix C Additional Experiment Results
-C.1 Experiment I
+Appendix E Additional Experiment Results
+E.1 Consistency in Inference Alignment
+We examine how closely our model’s generated responses align with a user-specified target objective, in this case, the 9th utility function (index 8).
+Figure 4 shows the distribution of normalized utility values across all ten reward models for the “harmless” vs.
+“humor” task, and Table 4 reports the corresponding mean normalized values.
+Notably, the 9th utility function (index 8) attains the highest average score of 0.68, demonstrating that the LLM effectively prioritizes the user-indicated objective.
+In other words, when prompts include the requirement <max_utility_index> i specifying the 9th utility, the model consistently generates responses that yield higher scores along that dimension than along the others.
+These findings confirm that our fine-tuned model can reliably follow the user-designated target objective during inference.
+Figure 4: Utility value distribution for the “harmless” vs. “humor” task. The required utility index is 8 (i.e., the 9th utility function).
+Table 4: Mean normalized value for each utility function.
+Utility index 0 1 2 3 4 5 6 7 8 9
+Mean 0.57 0.57 0.56 0.56 0.57 0.57 0.49 0.55 0.68 0.58
+E.2 Pareto Front Analysis: Four Other Tasks
 In addition to the main results presented in the paper, Figure 5 shows the Pareto fronts for the remaining four tasks in our experiments.
@@
 Figure 5: Pareto fronts of the remaining four tasks.
-C.2 Experiment III
-We investigate the impact of the online training stage by comparing three variants of our method: (1) pure offline training (w/o online), (2) one online iteration (w/ online iter 1), and (3) two online iterations (w/ online iter 2).
-As shown in Figure 8, each subplot corresponds to the Pareto fronts for a different task.
-Across all four tasks, w/o online (blue line) consistently produces the weakest Pareto front, indicating suboptimal performance when balancing multiple objectives.
-Introducing a single online iteration (w/ online iter 1, green line) noticeably expands the Pareto front, suggesting that even a brief online training phase can significantly improve multi-reward optimization.
-Finally, adding a second online iteration (w/ online iter 2, red line) further refines these trade-offs, achieving the most outward (and thus superior) Pareto front among the three variants.
-These results confirm the necessity of the online stage, demonstrating that multiple online iterations can better align the model with the target objectives.
-Figure 8: Ablation study of the online stage on the remaining four tasks.
-C.3 Experiment IV
-We examine how closely our model’s generated responses align with a user-specified target objective, in this case, the 9th utility function (index 8).
-Figure 9 shows the distribution of normalized utility values across all ten reward models for the “harmless” vs.
-“humor” task, and Table 6 reports the corresponding mean normalized values.
-Notably, the 9th utility function (index 8) attains the highest average score of 0.68, demonstrating that the LLM effectively prioritizes the user-indicated objective.
-In other words, when prompts include the requirement <max_utility_index> i specifying the 9th utility, the model consistently generates responses that yield higher scores along that dimension than along the others.
-These findings confirm that our fine-tuned model can reliably follow the user-designated target objective during inference.
-Figure 9: Utility value distribution for the “harmless” vs. “humor” task. The required utility index is 8 (i.e., the 9th utility function).
-Utility index 0 1 2 3 4 5 6 7 8 9
-Mean 0.57 0.57 0.56 0.56 0.57 0.57 0.49 0.55 0.68 0.58
-Table 6: Mean normalized value for each utility function.
-C.4 Experiment V
-We investigate how varying the number of utility functions (5, 10, 15) affects our method’s ability to balance multiple objectives.
-Figure 10 shows the resulting Pareto fronts across four different tasks, where an outward-shifted curve indicates better overall performance.
-We observe that using only 5 utility functions (blue line) yields a slightly weaker Pareto front compared to the other two settings.
-In contrast, the results for 10 (green line) and 15 (red line) utility functions are nearly identical, suggesting that increasing beyond 10 utility functions does not lead to a substantial performance gain.
-Consequently, our method demonstrates robustness to the exact number of utility functions, so long as it is sufficiently large to capture the relevant trade-offs among the objectives.
-Figure 10: Ablation study of different numbers of utility functions on remaining four tasks.
-C.5 Experiment VI
+E.3 Ablation Study
+We conduct an ablation study to investigate how other factors (e.g., alternative design choices, and hyperparameters) influence the performance of our method.
+Necessity of the online stage.
+We perform an ablation study to compare the performance of pure offline training (“w/o online”), a single online iteration (“w/ online iter 1”), and two online iterations (“w/ online iter 2”) across all tasks.
+We present the performance results for “harmless” vs.
+“humor” and “summary” vs.
+“faithful” tasks in Figure 6.
+We observe that models incorporating online iterations consistently outperform the offline-only baseline.
+Notably, even a single online iteration yields significant improvements, with further performance gains observed when two iterations are applied, thereby underscoring the necessity of the online stage in our training framework.
+The online stage is important because it enables the model to generalize over unseen samples in the offline dataset, and ultimately enhancing its ability to generate appropriate responses corresponding to the required utility function index.
+Figure 6: Ablation study of the online stage on all six tasks.
+Number of utility functions.
+We vary the number of utility functions from {5,10,15} and compare the resulting empirical Pareto fronts in Figure 7.
+With 5 utility functions, the generated Pareto front is slightly inferior to those produced with 10 and 15 utility functions.
+Specifically, the front with 5 utility functions captures a narrower range of optimal trade-offs across the objectives.
+However, when the number of utility functions is increased to 10 or 15, the resulting Pareto fronts exhibit comparable performance.
+Increasing the number of utility functions beyond 10 does not yield substantial gains.
+These results demonstrate that our method is insensitive to the precise number of utility functions, provided that a sufficient number (around 10) is used to effectively cover the preference space.
+In practical terms, employing 10 utility functions strikes a good balance between performance and computational efficiency, as further increases to 15 do not produce notable improvements.
+Figure 7: Ablation study of different numbers of utility functions on all six tasks.
+Variants of utility functions.
+We investigate an alternative design choice of the utility functions, i.e., linear functions.
+The linear function is designed in the form of $w^{T}x$ where $||w||_{2}=1$.
+To construct $N$ linear utility functions, we uniformly sample $N$ points from the unit circle and set $w$ accordingly.
+We plot the Pareto fronts using linear utility functions and our method across all tasks.
+The generated Pareto fronts, as shown in Figure 8, illustrate that our method continues to deliver strong results even with linear utility functions.
+However, there is still a performance gap between the results achieved with distributional Pareto-optimal utility functions and those obtained using linear functions.
+While linear functions still produce reasonably good trade-offs, they fail to capture the full range of optimal solutions attainable by our method, which leverages distributional utility representations.
+This highlights the advantage of our approach in handling more complex utility functions, demonstrating its superior flexibility in navigating multi-objective optimization.
 Figure 8: Abalation study of using linear utility functions.
-We compare the performance of our method when using linear utility functions (Ours (linear)) versus our original, distribution-based approach (Ours (param)) in Figure 11.
-In each subplot, the curves depict how effectively each method covers the trade-offs between two objectives.
-Although the linear approach maintains a reasonable Pareto front and thus remains effective for balancing multiple rewards, it consistently trails the distribution-based version in all tasks.
-This shortfall indicates that purely linear utility functions, while simpler to construct, capture a narrower range of potential trade-offs and consequently yield a less comprehensive frontier.
-In contrast, our distribution-based design is more flexible, enabling better alignment with diverse user preferences and resulting in a noticeably more outward Pareto front.
+Appendix F Proof
+The theorem in the main paper is restated here:
+Theorem 1 (Utility-Conditional Distributional Pareto-Optimality).
+Under Assumption 1, for any generated response $y\sim\pi_{\theta,i}(x)$ with reward vector $\mathbf{z}(y)$, let
+$i^{\star}=\arg\max_{j\in\{1,\dots,M\}}g_{\psi_{j}}(\mathbf{z}(y)).$
+Then the sub-policy $\pi_{\theta,i^{\star}}$ is distributionally Pareto-optimal.
+We first define the First-Order Stochastic Dominance (FOSD) for distributions and policies:
+Definition 2 (FOSD for distributions).
+Let $\boldsymbol{\mu}_{1},\boldsymbol{\mu}_{2}$ be probability measures on $\mathcal{Z}=[0,1]^{K}$.
+We say that $\boldsymbol{\mu}_{1}$ first-order stochastically dominates $\boldsymbol{\mu}_{2}$, written $\boldsymbol{\mu}_{1}\succ_{\mathrm{SD}}\boldsymbol{\mu}_{2}$, iff
+$\forall\,g_{\psi_{i}}\in\mathcal{U}^{\uparrow},\quad\mathbb{E}_{\mathbf{z}\sim\boldsymbol{\mu}_{1}}[g_{\psi_{i}}(\mathbf{z})]\geq\mathbb{E}_{\mathbf{z}\sim\boldsymbol{\mu}_{2}}[g_{\psi_{i}}(\mathbf{z})].$
+Definition 3 (FOSD for policies).
+For any policy $\pi$, let $\boldsymbol{\mu}(\pi)$ denote the return distribution induced by $\pi$.
+Policies $\pi_{1}$ and $\pi_{2}$ are ordered by FOSD according to Definition 2: $\pi_{1}\succ_{\mathrm{SD}}\pi_{2}$ iff $\boldsymbol{\mu}(\pi_{1})\succ_{\mathrm{SD}}\boldsymbol{\mu}(\pi_{2})$.
+With the definition of FOSD, we have the following lemma:
+Lemma 1 (Expected-utility separation under FOSD).
+Let $u:\mathcal{Z}\to\mathbb{R}$ be strictly increasing in each coordinate. If $\boldsymbol{\mu}_{1}\succ_{\mathrm{SD}}\boldsymbol{\mu}_{2}$ then
+$\mathbb{E}_{\mathbf{z}\sim\boldsymbol{\mu}_{1}}[u(\mathbf{z})]>\mathbb{E}_{\mathbf{z}\sim\boldsymbol{\mu}_{2}}[u(\mathbf{z})].$
+Proof.
+By Definition 3 the expected utilities are weakly ordered for all non-decreasing $u$ and strictly ordered for at least one.
+Because $u$ is strictly increasing, it belongs to the latter subset, hence the inequality is strict.
+∎
+Now we provide the detailed proof of Theorem 1:
+Proof.
+We aim to show that the sub-policy $\pi_{\theta,i^{\star}}$ is distributionally Pareto-optimal.
+By definition, a policy is distributionally Pareto-optimal if no other policy stochastically dominates it.
+Assume, for the sake of contradiction, that there exists another policy $\pi^{\prime}$ such that $\pi^{\prime}$ stochastically dominates $\pi_{\theta,i^{\star}}$, i.e., $\pi^{\prime}\succ_{\mathrm{SD}}\pi_{\theta,i^{\star}}$.
+According to Definition 3, this implies that the return distribution of $\pi^{\prime}$, denoted by $\boldsymbol{\mu}(\pi^{\prime})$, first-order stochastically dominates the return distribution of $\pi_{\theta,i^{\star}}$, denoted by $\boldsymbol{\mu}(\pi_{\theta,i^{\star}})$:
+$\boldsymbol{\mu}(\pi^{\prime})\succ_{\mathrm{SD}}\boldsymbol{\mu}(\pi_{\theta,i^{\star}}).$
+By Definition 2, this means that for all non-decreasing utility functions $u\in\mathcal{U}^{\uparrow}$, we have:
+$\mathbb{E}_{\mathbf{z}\sim\boldsymbol{\mu}(\pi^{\prime})}[u(\mathbf{z})]\geq\mathbb{E}_{\mathbf{z}\sim\boldsymbol{\mu}(\pi_{\theta,i^{\star}})}[u(\mathbf{z})],$
+and the inequality is strict for at least one such utility function.
+Now, consider the specific utility function $g_{\psi_{i^{\star}}}$.
+From Section 3.1 of the main paper, we know that $g_{\psi_{i^{\star}}}$ is strictly increasing in each coordinate (due to the added linear term with $\epsilon>0$).
+Therefore, we can apply Lemma 1 with $u=g_{\psi_{i^{\star}}}$, $\boldsymbol{\mu}_{1}=\boldsymbol{\mu}(\pi^{\prime})$, and $\boldsymbol{\mu}_{2}=\boldsymbol{\mu}(\pi_{\theta,i^{\star}})$.
+This gives us:
+$\mathbb{E}_{\mathbf{z}\sim\boldsymbol{\mu}(\pi^{\prime})}[g_{\psi_{i^{\star}}}(\mathbf{z})]>\mathbb{E}_{\mathbf{z}\sim\boldsymbol{\mu}(\pi_{\theta,i^{\star}})}[g_{\psi_{i^{\star}}}(\mathbf{z})].\quad(*)$
+However, by Assumption 1, the sub-policy $\pi_{\theta,i^{\star}}$ is defined as a policy that maximizes the expected value of the utility function $g_{\psi_{i^{\star}}}$.
+This means that for any other policy $\pi^{\prime}$, including the one we assumed to stochastically dominate $\pi_{\theta,i^{\star}}$, the following must hold:
+$\mathbb{E}_{\mathbf{z}\sim\boldsymbol{\mu}(\pi_{\theta,i^{\star}})}[g_{\psi_{i^{\star}}}(\mathbf{z})]\geq\mathbb{E}_{\mathbf{z}\sim\boldsymbol{\mu}(\pi^{\prime})}[g_{\psi_{i^{\star}}}(\mathbf{z})].$
+This inequality directly contradicts the strict inequality we derived from the assumption that $\pi^{\prime}$ stochastically dominates $\pi_{\theta,i^{\star}}$.
+Therefore, our initial assumption that such a policy $\pi^{\prime}$ exists must be false.
+Since no other policy stochastically dominates $\pi_{\theta,i^{\star}}$, by the definition of distributional Pareto-optimality, the sub-policy $\pi_{\theta,i^{\star}}$ is indeed distributionally Pareto-optimal.
+∎
+Appendix G User Study
+To evaluate the qualitative performance of UC-MOA compared to the state-of-the-art method RiC, we conducted a user study where participants rated anonymized responses from both systems to a series of diverse prompts.
+This section details the design of the study, the presentation of prompts and system responses, and an analysis of the collected user preferences and qualitative feedback.
+G.1 User Study Design
+Objective: The primary goal of this study was to compare user preferences for responses generated by UC-MOA versus RiC across a range of tasks requiring a blend of humor, information, and practicality.
+Participants: We utilized the criterion sampling method [24] to recruit participants.
+The participants were expected to have English reading skills and hold some basic knowledge of large language models.
+Task: For each of the five prompts, participants were presented with two anonymized system responses.
+The assignment of responses to "System A" and "System B" was randomly permuted for each prompt and each participant to prevent bias.
+Participants were not informed which system corresponded to UC-MOA or RiC.
+Participants were asked to choose which of the two responses they preferred and were given the option to provide a brief written justification for their choice.
+G.1.1 Participant Instructions
+We first present the following general instructions:
+Welcome to our study on evaluating AI-generated text.
+Your task is to help us understand which responses people find more humor.
+We will present 5 questions below.
+For each question, you will see a prompt that was given to two different AI systems.
+Below the prompt, you will see two responses, labeled “System A” and “System B”.
+These labels are randomized.
+Please read both responses carefully and select the response (System A or System B) that you prefer.
+Optionally, you can provide a brief explanation for your choice in the text box provided.
+Your feedback is valuable in helping us understand your preferences.
+There are no right or wrong answers; we are interested in your genuine preference.
+Thank you for your participation!
+Here are five prompts and the corresponding responses:
+  1. 1.
+    Explain quantum computing to a 12–year–old in a funny way.
+    • System A: Quantum computing is like giving your homework to a coin that can be both heads and tails at once—and the coin roasts you about Schrödinger’s cat while it solves algebra!
+    • System B: Imagine your game console pressing every button at the same time to find the winning combo, all while cracking puns about super positions.
+  2. 2.
+    I’m anxious about my job interview tomorrow. Give me three concrete tips that also lighten the mood.
+    • System A: 1. Breathe like Darth Vader on a beach holiday—slooow in, slooow out. 2. Write your top strengths on a sticky note and fistbump it like a tiny hypebuddy.
+    • System B: Picture your nerves as a mischievous house-cat: acknowledge it, pet it once, then move it off your keyboard so you can prep.
+  3. 3.
+    Write a short, humorous yet informative comparison between cats and dogs as pets.
+    • System A: Cats are the introverted artists of the animal kingdom: lowmaintenance, occasionally aloof, and prone to 3 a.m.
+parkour.
+Dogs are the extroverted linebackers: loyalty dialed up to 11, endless fetch, and zero respect for personal space.
+    • System B: Dog owners get free cardio sessions (walkies!), while cat owners get free zen lessons (meditative ignoring). Both come with bonus fur-glitter for your clothes.
+  4. 4.
+    Help me troubleshoot why my sourdough starter isn’t rising, with a dash of comedy.
+    • System A: If your starter is lazier than a teenager on a Sunday, first make sure it’s actually getting food: equalweight flour and water, twice a day.
+Think of feeding it as paying rent to billions of yeasty tenants.
+    • System B: Pretend your starter is a micro-pet dragon: it needs warmth (25 extdegreeC), snacks (carbs), and breathing room. Starve it or freeze it and you’ll have a very sleepy dragon.
+  5. 5.
+    Draft a witty but practical email reminding my roommate to pay their share of the rent.
+    • System A: Subject: Friendly Rent Reminder
+      Hey [Roomie], just popping into your inbox like the firstofthemonth fairy. The rent goblin is hungry for your share (\$650) by Friday.
+    • System B: Subject: The First of the Month Strikes Again
+      Dearest co-dweller, our beloved landlord prefers money to memes, and the ritual tribute (\$650) falls due this week.
+All system outputs are truncated to the first two sentences for brevity.
+G.2 Representative Participant Comments
+Participants were invited (optionally) to write a brief justification after each vote. Below are representative, lightly edited remarks that capture common themes.
+  • P1: System A explained the idea more clearly while still making me laugh—System B’s joke landed but felt vague.
+  • P2: A’s breathing tip and sticky-note trick were immediately actionable; B’s cat metaphor was cute but less useful.
+  • P3: I liked A’s vivid imagery of 3 a.m. parkour; B was funny too but didn’t contrast the pets as sharply.
+  • P4: System B’s dragon analogy was hilarious, but A actually told me what temperature to use—so I picked A.
+  • P5: Both emails were witty, but A gave a concrete amount and deadline, which felt more practical.
+These qualitative comments mirror the quantitative outcome: participants consistently favored answers that blended humor with concrete, actionable advice—qualities more frequently exhibited by UC-MOA outputs.
+G.3 Overall Findings and Discussion
+The user study results, encompassing both the quantitative preferences (46 out of 60 favor our method while only 14 out of 60 favor RiC) and the qualitative participants’ comments, reveal a consistent pattern.
+Participants generally favored responses that effectively balanced humor with clear, concrete, and actionable information.
+The qualitative feedback, as exemplified by the comments in Section G.2, indicates that while both systems were capable of generating humorous or engaging text, the perceived utility and directness of the advice often served as key differentiators in user preference.
+These attributes were more consistently observed in the outputs generated by UC-MOA, contributing to its higher overall preference rate in this study.
+This suggests that UC-MOA demonstrates a stronger ability to integrate an engaging style with practical substance, thereby aligning more closely with user expectations for the types of prompts investigated.